### v1

# references.bib

Abstract
Rank-based statistical metrics, such as the invariant statistical loss (ISL), have recently emerged as robust and practically effective tools for training implicit generative models. In this work, we introduce dual-ISL, a novel likelihood-free objective for training implicit generative models that interchanges the roles of the target and model distributions in the ISL framework, yielding a convex optimization problem in the space of model densities. We prove that the resulting rank-based discrepancy $d_{K}$ is i) continuous under weak convergence and with respect to the $L^{1}$ norm, and ii) convex in its first argument—properties not shared by classical divergences such as KL or Wasserstein distances. Building on this, we develop a theoretical framework that interprets $d_{K}$ as an $L^{2}$-projection of the density ratio $q=p/\tilde{p}$ onto a Bernstein polynomial basis, from which we derive exact bounds on the truncation error, precise convergence rates, and a closed-form expression for the truncated density approximation. We further extend our analysis to the multivariate setting via random one-dimensional projections, defining a sliced dual-ISL divergence that retains both convexity and continuity. We empirically show that these theoretical advantages translate into practical ones. Specifically, across several benchmarks dual-ISL converges more rapidly, delivers markedly smoother and more stable training, and more effectively prevents mode collapse than classical ISL and other leading implicit generative methods—while also providing an explicit density approximation.
1 Introduction
Implicit generative models are a class of models that learn to generate data samples without explicitly modeling the underlying probability distribution [mohamed2016learning], enabling flexible modeling of high-dimensional data across vision (e.g., DCGAN [radford2015unsupervised]), audio (e.g., WaveGAN [donahue2018adversarial]), and text domains (e.g., SeqGAN [yu2017sequence]). Instead of directly estimating the data distribution, these models learn a mapping from a simple input distribution (such as a multivariate Gaussian) to the data space through a deterministic or stochastic function. A prominent example is the generator in a Generative Adversarial Network (GAN) [goodfellow2014generative], which transforms random noise vectors into realistic data samples. The generator is trained in tandem with a discriminator that learns to distinguish real data from generated data, providing feedback that guides the generator to improve. Unlike explicit models, implicit models do not require tractable likelihoods, allowing them to generate high-quality samples even when the data distribution is complex or high-dimensional.
The Invariant Statistical Loss (ISL) is a rank-based loss function recently proposed in [de2024training] that compares the empirical order statistics of samples from the data and from the implicit generative model. In this work, we introduce dual-ISL, a novel likelihood-free objective obtained by swapping the roles of the data and model distributions within the ISL framework. Remarkably, the induced discrepancy $d_{K}$ admits a fully explicit closed-form density approximation: it is exactly the $L^{2}$–projection of the density ratio
$q(x)\;=\;\frac{p_{\mathrm{target}}(x)}{p_{\mathrm{model}}(x)},\qquad x=F_{\mathrm{target}}^{-1}(t),\;t\in[0,1],$
onto the space of dual-Bernstein polynomials of degree $K$ [lorentz2012bernstein, jiittler1998dual]. Writing
$q_{K}(x)=\sum_{n=0}^{K}\mathbb{Q}_{K}(n)\,\tilde{b}_{n,K}\bigl{(}F_{\mathrm{target}}(x)\bigr{)}$
with computable coefficients $\{\mathbb{Q}_{K}(n)\}$ immediately yields
$p_{\mathrm{model}}(x)\;\approx\;\frac{p_{\mathrm{target}}(x)}{q_{K}(x)}.$
This explicit representation not only provides analytic error bounds via Bernstein approximation theory and ensures convexity over the space of densities, but also enables efficient density evaluation without auxiliary sampling and provable convergence rates inherited from polynomial approximation. By marrying likelihood-free training with a tractable, closed-form density, dual-ISL bridges a critical gap—offering both rigorous theory and practical stability in implicit generative modeling. Moreover, this rank-based construction directly parallels the univariate optimal transport problem: matching order statistics via the probability–integral transform recovers the Monge map and yields the $p$-Wasserstein distance [villani2008optimal]. Dual-ISL extends this perspective by providing an explicit, closed-form polynomial approximation of the density ratio—complete with convexity guarantees and convergence rates—rather than only a transport map.
2 The Invariant Statistical Loss (ISL)
We briefly review the invariant statistical loss (ISL) from [de2024training]. ISL is built on a simple rank statistic whose distribution is exactly uniform when two samples come from the same probability density function, and which varies continuously under $L^{1}$-perturbations of the underlying densities.
2.1 Rank statistic and uniformity
Let $\tilde{y}_{1},\dots,\tilde{y}_{K}$ be i.i.d. samples from a univariate real distribution with pdf $\tilde{p}$, and let $y$ be a single sample independently drawn from another distribution with pdf $p$. Define the subset
$\mathcal{A}_{K}\;:=\;\Bigl{\{}\tilde{y}\in\{\tilde{y}_{k}\}_{k=1}^{K}:\tilde{y}\leq y\Bigr{\}},$
and the rank statistic
$A_{K}\;:=\;\bigl{|}\mathcal{A}_{K}\bigr{|},$ (1)
i.e., $A_{K}$ counts how many samples in $\{\tilde{y}_{1},\ldots,\tilde{y}_{K}\}$ lie at or below $y$. Then $A_{K}$ is a discrete random variable (r.v.) taking values in $\{0,1,\ldots,K\}$, and we denote its pmf by
$\mathbb{Q}_{K}:\{0,\ldots,K\}\to[0,1].$
When the two pdfs $p$ and $\tilde{p}$ coincide, this pmf is exactly uniform [de2024training].
Theorem 2.1.
If $p=\tilde{p}$, then $A_{K}$ is uniformly distributed on $\{0,\ldots,K\}$, i.e. $\mathbb{Q}_{K}(n)=\tfrac{1}{K+1}$ for all $n\in\{0,\ldots,K\}$.
2.2 ISL discrepancy
The ISL discrepancy quantifies the deviation of the pmf $\mathbb{Q}_{K}$ from the uniform law on $\{0,\dots,K\}$. To be specific, we define the discrepancy function
$d_{K}(p,\tilde{p}):=\dfrac{1}{K+1}\|\mathbb{Q}_{K}-\mathbb{U}_{K}\|_{\ell_{1}}=\frac{1}{K+1}\sum_{n=0}^{K}\Bigl{|}\,\tfrac{1}{K+1}\;-\;\mathbb{Q}_{K}(n)\Bigr{|}=\dfrac{2}{K+1}\mathrm{TV}(\mathbb{Q}_{K},\mathbb{U}).$ (2)
where $\mathbb{U}_{K}$ is the uniform pmf on $\{0,\dots,K\}$ and $\mathrm{TV}(\cdot,\cdot)$ denotes total variation distance. By Theorem 2.1, $d_{K}(p,p)=0$ for all $K$. Moreover, Theorem 2.2 below, ensures that $d_{K}(p,\tilde{p})$ depends continuously on $\tilde{p}$ in the $L^{1}$ sense, while Theorem 2.3 guarantees that if $d_{K}(p,\tilde{p})=0$ for all $K$, then $\tilde{p}=p$ almost everywhere. Hence, in the large-$K$ limit, $d_{K}$ behaves as a proper divergence, vanishing precisely when the two densities coincide.
Theorem 2.2 (Continuity).
If $\|p-\tilde{p}\|_{L^{1}(\mathbb{R})}\,\leq\,\epsilon,$ then for all $n\in\{0,\ldots,K\}$,
$\frac{1}{K+1}-\epsilon\;\;\leq\;\;\mathbb{Q}_{K}(n)\;\;\leq\;\;\frac{1}{K+1}+\epsilon.$
Theorem 2.3 (Identifiability).
Let $p,\tilde{p}$ be pdfs of univariate real r.v.s. If the rank statistic $A_{K}$ in (1) is uniformly distributed on $\{0,\ldots,K\}$ for every $K\in\mathbb{N}$, then $p=\tilde{p}$ almost everywhere.
Finally, when $\tilde{p}=\tilde{p}_{\theta}$ depends smoothly on a parameter vector  $\theta$, one can show (under mild regularity assumptions) that $\theta\mapsto d_{K}(p,\tilde{p}_{\theta})$ is continuous and differentiable, making it suitable for gradient-based optimization (see Theorem 4 in [de2024robust]). For full proofs and additional remarks, see [de2024training, de2024robust].
2.3 A surrogate for ISL optimization
Directly minimizing the divergence $d_{K}\bigl{(}p,\tilde{p}_{\theta}\bigr{)}$ with respect to the generator parameters $\theta$ is normally not feasible: the pmf $\mathbb{Q}_{K}$ has to be approximated empirically and its dependence on $\theta$ is unknown. To overcome this difficulty, [de2024training] introduced a carefully designed surrogate loss that (i) closely tracks $d_{K}$, and (ii) admits gradient optimization via standard backpropagation. This surrogate is constructing by approximating the pmf of $\mathbb{Q}_{K}$ using sigmoidal functions and a Gaussian kernel density estimator. In practice, training with the surrogate yields virtually identical performance to optimizing the true ISL, while remaining fully likelihood-free and amenable to efficient stochastic optimization. For full details of the surrogate derivation, implementation, and bias-variance trade-offs, see [de2024robust, Section 2.3].
3 The dual-Invariant statistical loss
By interchanging the roles of the data distribution $p$ and the model distribution $\tilde{p}$ in the ISL framework, we obtain a dual objective that remains likelihood-free, but crucially becomes convex in the model pdf $\tilde{p}$.
3.1 Continuity and convexity of $d_{K}(p,\tilde{p})$
[…]
A key insight in the continuity proof is that each probability mass $\mathbb{Q}_{K}(n)$, for $n=0,\ldots,K$, can be written as a continuous mixture of the binomial pmf’s. Indeed, drawing $K$ i.i.d. samples $\tilde{y}_{i}\sim\tilde{p}$ and counting how many of them fall below $y$ yields a $\mathrm{Binomial}(K,\tilde{F}(y))$ distribution. Since $y$ itself is drawn from $p$, one obtains [de2024training, Appendix 1]
$\mathbb{Q}_{K}(n)$ $:=\int_{\mathbb{R}}h_{n}(y)\,p(y)\,\mathrm{d}y\,,$
where, $\;h_{n}(y):=\binom{K}{n}\,\tilde{F}(y)^{n}\bigl{(}1-\tilde{F}(y)\bigr{)}^{K-n}\,,\quad n=0,1,\dots,K\,.$
and the bounded, continuous functions $h_{n}$ then ensure weak continuity of $\mathbb{Q}_{K}$ and hence of $d_{K}$. We formalize this argument in the following theorem.
Theorem 3.1 (Continuity under weak convergence).
Let $(p_{n})_{n\geq 1}$ be a sequence of pdfs on $\mathbb{R}$ converging weakly to a density $p$, and let $\tilde{p}$ be a fixed reference density with cdf $\tilde{F}$. For each $K\in\mathbb{N}$, define
$\mathbb{Q}_{K}^{(n)}(m):=\int_{\mathbb{R}}\binom{K}{m}\tilde{F}(y)^{m}\bigl{(}1-\tilde{F}(y)\bigr{)}^{K-m}\,p_{n}(y)\,\mathrm{d}y,$
for $m=0,\dots,K$. Then
  1. (i)
    (Pointwise convergence) $\;\lim_{n\to\infty}\mathbb{Q}_{K}^{(n)}(m)\;=\;\mathbb{Q}_{K}(m)\;$ for each $\;m=0,\dots,K$.
  2. (ii)
    (Continuity of $d_{K}$) $\;\lim_{n\to\infty}d_{K}(p_{n},\tilde{p})\;=\;d_{K}(p,\tilde{p})\,$.
Proof.
See Appendix A. ∎
Since strong convergence implies weak convergence, the previous theorem remains applicable when the sequence $\{p_{n}\}_{n\geq 1}$ converges to $p$ in the $L^{1}$ norm. We can also establish that the ISL divergence is continuous with respect to its second argument $\tilde{p}$; a detailed proof can be found in Appendix A, Theorem A.1.
We now see that the discrepancy $d_{K}(p,\tilde{p})$ is indeed convex in its first argument.
Theorem 3.2 (Convexity).
For any probability distributions $p_{1},p_{2}$ and $\tilde{p}$ on $\mathbb{R}$, and for any $\lambda\in[0,1]$, the discrepancy $d_{K}$ satisfies
$d_{K}(\lambda p_{1}+(1-\lambda)p_{2},\,\tilde{p})\leq\lambda\,d_{K}(p_{1},\,\tilde{p})+(1-\lambda)\,d_{K}(p_{2},\,\tilde{p})$
Proof.
See Appendix A. ∎
3.2 A dual loss function
Because $d_{K}$ is convex in its first argument, we can obtain a new training criterion by swapping the data and model distributions. Specifically, let $\tilde{y}\sim\tilde{p}$ be a simulated sample from our generator and let $y_{1:K}\overset{\mathrm{i.i.d.}}{\sim}p$ be $K$ independent real data points. We then form the rank statistic as
$\tilde{A}_{K}\;:=\;\left\lvert\Bigl{\{}y\in\{y_{k}\}_{k=1}^{K}:y\leq\tilde{y}\Bigr{\}}\right\rvert$
whose pmf ~$\tilde{\mathbb{Q}}_{K}(n)=\mathbb{P}\left(\tilde{A}_{K}=n\right)$ remains uniform if and only if $\tilde{p}=p$. All of our previous guarantees—continuity under small $L^{1}$ perturbations (Theorem 2.2) and identifiability when $\mathbb{Q}_{K}$ is exactly uniform for every $K$ (Theorem 2.3)—carry over unchanged. The dual-ISL discrepancy
$d_{K}(\tilde{p},p)=\frac{1}{K+1}\sum_{n=0}^{K}\left|\tilde{\mathbb{Q}}_{K}(n)-\dfrac{1}{K+1}\right|$
therefore yields a convex, likelihood-free training objective in the space of generator densities. The pseudocode for this method is provided in the supplementary material (see Algorithm 1).
3.3 Dual-ISL vs. ISL, GANs & Diffusion on 1D distributions
We start considering the same experimental setup as [zaheer2017gan, de2024training]. We evaluate dual-ISL on six benchmark targets using $N=1000$ i.i.d. samples drawn from each distribution. The first three are standard univariate pdfs, and the latter three are mixtures with equal mixing weights. Model1 combines Gaussians $\mathcal{N}(5,2)$ and $\mathcal{N}(-1,1)$; Model2 combines Gaussians $\mathcal{N}(5,2)$, $\mathcal{N}(-1,1)$, and $\mathcal{N}(-10,3)$; and Model3 combines a Gaussian $\mathcal{N}(-5,2)$ with a Pareto$(5,1)$ distribution.
We train a 4-layer MLP generator (7–13–7–1 units, ELU activations) with $\epsilon\sim\mathcal{N}(0,1)$ input noise for $10^{4}$ epochs with Adam (learning rate $10^{-2}$), and compare Dual-ISL, ISL, GAN [goodfellow2014generative], WGAN [arjovsky2017wasserstein], MMD-GAN [li2017mmd], and a diffusion baseline using the Kolmogorov–Smirnov distance (KSD) metric (Table 1). Experimental details are provided in Supplementary Material Section D.1.
Target Dual-ISL ISL GAN WGAN MMD-GAN Diffusion
$\mathcal{N}(4,2)$ $\mathbf{0.018\pm 0.005}$ $0.020\pm 0.003$ $0.018\pm 0.003$ $0.024\pm 0.017$ $0.042\pm 0.026$ $0.020\pm 0.002$
$\mathcal{U}(-2,2)$ $0.034\pm 0.015$ $0.021\pm 0.004$ $0.049\pm 0.032$ $0.064\pm 0.062$ $0.104\pm 0.060$ $\mathbf{0.013\pm 0.002}$
$\mathrm{Cauchy}(1,2)$ $0.016\pm 0.003$ $\mathbf{0.013\pm 0.002}$ $0.013\pm 0.002$ $0.052\pm 0.055$ $0.031\pm 0.008$ $0.114\pm 0.034$
$\mathrm{Pareto}(1,1)$ $\mathbf{0.090\pm 0.080}$ $0.198\pm 0.148$ $0.117\pm 0.041$ $0.106\pm 0.043$ $0.158\pm 0.168$ $0.209\pm 0.011$
$\mathrm{Mixture}_{1}$ $0.016\pm 0.004$ $\mathbf{0.016\pm 0.002}$ $0.017\pm 0.004$ $0.080\pm 0.069$ $0.054\pm 0.033$ $0.031\pm 0.031$
$\mathrm{Mixture}_{2}$ $\mathbf{0.016\pm 0.002}$ $0.017\pm 0.003$ $0.026\pm 0.014$ $0.031\pm 0.023$ $0.042\pm 0.061$ $0.050\pm 0.005$
$\mathrm{Mixture}_{3}$ $\mathbf{0.170\pm 0.019}$ $0.171\pm 0.012$ $0.190\pm 0.094$ $0.216\pm 0.040$ $0.187\pm 0.108$ $0.173\pm 0.024$
Table 1: KSD over 10 runs for Dual-ISL and baselines. Setup: $K=10$, 1000 epochs, $N=1000$.
The convexity of the dual-ISL objective not only accelerates convergence—yielding faster, smoother, and more stable training curves compared to classical ISL (see Figure 7 in Appendix)—but also enhances mode coverage on challenging mixtures. As shown in Figure 1, both dual-ISL and classical ISL successfully avoid the mode collapse exhibited by MMD-GAN, with dual-ISL most accurately capturing the heavy tail of the Pareto component.
(a) dual-ISL
(b) ISL
(c) MMD-GAN
Figure 1: Comparison of dual-ISL, standard ISL, and MMD-GAN for modeling a mixture of Pareto and Normal distributions. Subfigure 1(a) displays the dual-ISL results, Subfigure 1(b) illustrates the performance of the standard ISL approach, and Subfigure 1(c) showcases the outcomes obtained via MMD-GAN. Further comparisons—including diffusion models and additional target distributions—are provided in Appendix D
Additional experiments in the supplementary material provide detailed runtime benchmarks, demonstrating the computational advantages of dual-ISL over the standard ISL formulation. In Appendix D.3, we also propose a new ISL-based method with a monotonicity penalty that guarantees recovery of the optimal-transport map even for distributions without finite moments (e.g., heavy-tailed), an advantage over the $p$-Wasserstein distance which requires finite $p$th moments.
4 An $L^{2}$-projection view of $d_{K}$
We adopt a projection-based view of ISL. From this point on, we treat $p$ and $\tilde{p}$ interchangeably—so that, with a slight abuse of notation, our framework covers both standard ISL ($q=p/\tilde{p}$) and dual-ISL ($q=\tilde{p}/p$). Specifically, we show that the discrete pmf $\mathbb{Q}_{K}(n)$ coincides with the $L^{2}$-projection coefficients of the density ratio $q=p(x)/\tilde{p}(x)$ onto the degree-$K$ Bernstein basis $\{b_{n,K}\}_{n=0}^{K}$ In this light, ISL becomes a purely likelihood-free density-ratio divergence—comparing projection coefficients rather than intractable likelihoods —and we conclude by deriving sharp convergence rate bounds.
4.1 Projection interpretation
To reveal the underlying geometry of $d_{K}(\cdot,\cdot)$, we define a linear operator that collects the $K+1$ probabilities $\mathbb{Q}_{K}(0),\ldots,\mathbb{Q}_{K}(K)$, into a single vector. We then show that each entry $\mathbb{Q}_{K}(n)$ is precisely the $L^{2}$ inner product between a density ratio and its corresponding Bernstein basis function.
Definition 4.1 (Binomial mapping).
Let $p,\tilde{p}\in C(\mathbb{R})$ be two continuous pdfs with cdfs $F$ and $\tilde{F}$. For any integer $K\geq 1$, define the operator,
$\Phi_{K}(p,\tilde{p}):=\bigl{(}\mathbb{Q}_{K}(0),\,\mathbb{Q}_{K}(1),\,\dots,\,\mathbb{Q}_{K}(K)\bigr{)}\;\in\;\mathbb{R}^{K+1}.$
It is straightforward from the integral representation that, for each fixed $\tilde{p}$, the map $p\;\mapsto\;\Phi_{K}(p,\tilde{p})$ is linear and continuous under mild regularity conditions on $p$ and $\tilde{p}$. A full statement of these and related properties appears in Theorem B.1.
The next result shows that $\Phi_{K}$ admits a Riesz representation (see [brezis2011functional, Theorem 4.11]), expressing each probability mass $\mathbb{Q}_{K}$ as an $L^{2}$ inner product with a Bernstein basis function. Let the $n$th Bernstein polynomial of degree $K$ be defined as
$b_{n,K}(t)\;:=\;\binom{K}{n}\,t^{n}\,(1-t)^{K-n},\qquad t\in[0,1].$
Theorem 4.1.
(Riesz representation of $\Phi_{K}$) Let $\tilde{p}$ be a fixed continuous density on $\mathbb{R}$ with cdf $\tilde{F}$. Then for any $K\geq 0$, the operator $\Phi_{K}(\cdot,\tilde{p})$ mapping $p\mapsto(\mathbb{Q}_{K}(0),\dots,\mathbb{Q}_{K}(K))$ satisfies
$\mathbb{Q}_{K}(n)=\int_{\mathbb{R}}b_{n,K}\bigl{(}\tilde{F}(x)\bigr{)}\,p(x)\,\mathrm{d}x=\bigl{\langle}\,b_{n,K}\circ\tilde{F},\;p\bigr{\rangle}_{L^{2}(\mathbb{R})},\qquad n=0,\dots,K.$
Moreover, if $\tilde{p}(x)>0$ for all $x\in\mathbb{R}$, then defining the density ratio $q(x)=p(x)/\tilde{p}(x)$, we get
$\mathbb{Q}_{K}(n)=\int_{0}^{1}b_{n,K}(t)\,q\bigl{(}\tilde{F}^{-1}(t)\bigr{)}\,\mathrm{d}t=\bigl{\langle}\,b_{n,K},\,q\circ\tilde{F}^{-1}\bigr{\rangle}_{L^{2}([0,1])},\quad n=0,\ldots,K.$
Proof.
See Appendix B. ∎
Theorem 4.1 implies that, if we define $\tilde{q}(t)=q\bigl{(}\tilde{F}^{-1}(t)\bigr{)}$ for $t\in[0,1]$ then each coefficient $\mathbb{Q}_{K}(n)$ is exactly the $L^{2}([0,1])$-projection of $\tilde{q}$ onto the Bernstein polynomial $b_{n,K}$. Equivalently, the vector $\{\mathbb{Q}_{K}(n)\}_{n=0}^{K}$ collects the best mean-square approximation coefficients of the push-forward density ratio $q\circ\tilde{F}^{-1}$ in the degree-$K$ Bernstein basis.
Theorem 4.2 (Bernstein-basis truncation for the density-ratio).
Let $p,\tilde{p}\in C(\mathbb{R})$ with $\tilde{p}(x)>0$ for all $x\in\mathbb{R}$. Then $\tilde{q}\in C([0,1])$ admits the Bernstein-polynomial expansion
$\tilde{q}(t)\;=\;\sum_{n=0}^{\infty}\alpha_{n}\,b_{n,K}(t),$
where $\alpha=(\alpha_{n})_{n\geq 0}$ are the unique Bernstein-basis coordinates of $\tilde{q}$. Its degree-$K$ truncation can be expressed as
$\tilde{q}_{K}(t)\;:=\;\sum_{n=0}^{K}\alpha_{n}\,b_{n,K}(t)=\sum_{n=0}^{K}\mathbb{Q}_{K}(n)\,\tilde{b}_{n,K}(t),$
where $\{\tilde{b}_{n,K}\}_{n=0}^{K}$ is the dual Bernstein basis [jiittler1998dual].
Proof.
See Proof in Appendix B. ∎
Remark 4.1.
With $\tilde{p}$ fixed, the map $\Phi_{K}:p\;\longmapsto\;\bigl{(}\mathbb{Q}_{K}(0),\dots,\mathbb{Q}_{K}(K)\bigr{)}$, cannot distinguish between any two target densities $p_{1},p_{2}$ whose pushed-through ratios $q_{i}\circ\tilde{F}^{-1}$ have the same degree-$K$ Bernstein projections. Equivalently,
$\Phi_{K}(p_{1})=\Phi_{K}(p_{2})\quad\Longleftrightarrow\quad\bigl{\langle}b_{n,K},\,q_{1}\circ\tilde{F}^{-1}-q_{2}\circ\tilde{F}^{-1}\bigr{\rangle}_{L^{2}([0,1])}=0\quad\forall\,n=0,\dots,K.$
Thus $\Phi_{K}$ factors through the quotient of $L^{2}([0,1])$ by the subspace orthogonal to $\mathrm{span}\{b_{n,K}\}$, inducing a bijection onto its image.
4.2 Approximation error
We now quantify the truncation error in the approximation ratio $\tilde{q}_{K}(t)$ is an estimate of $q(x)$ and it remains uniformly close to 1, with its sup-norm deviation bounded by the discrepancy $d_{K}(p,\tilde{p})$.
Theorem 4.3.
Let $p,\tilde{p}\in C(\mathbb{R})$ be pdfs. Then $q_{K}(x)$ satisfies
$\|q_{K}-1\|_{\infty}\leq(K+1)^{2}\,d_{K}(p,\tilde{p}).$
Proof.
See Proof in Appendix B
By standard Bernstein-approximation theory, one can bound the truncation error via the modulus of continuity of $q$. In particular, if $q\in C^{2}(\mathbb{R})$, then
$\|q-q_{K}\|_{\infty}=O(K^{-1}),$
and more generally, if $q$ is $\alpha$–Hölder continuous on $\mathbb{R}$, then
$\|q-q_{K}\|_{\infty}=O\bigl{(}K^{-\alpha/2}\bigr{)}.$
See [gzyl1997weierstrass] for the $C^{2}$ case and [mathe1999approximation] for the Hölder regime.
Remark 4.2.
If we assume that $q\in C^{2}(\mathbb{R})$, by the triangle inequality, we have
$\|q(x)-1\|_{\infty}\leq\|1-q_{K}(x)\|_{\infty}+\|q_{K}(x)-q(x)\|_{\infty}\leq(K+1)^{2}d_{K}(p,\tilde{p})+\dfrac{\|q(x)^{\prime\prime}\|_{\infty}}{8K}.$ (3)
To empirically validate Equation 3, we train the same NN architecture under identical hyperparameters as in our earlier experiments. The model receives as input noise $z\sim\mathcal{N}(0,1)$ and approximates a mixture of Cauchy distributions. We then recover the estimated density $\tilde{p}$ via kernel density estimation and compute the second derivative of the quotient $q$ with sixth-order central finite differences. Each experiment is repeated ten times, and the mean results are plotted in Figure 2.
(a) Fixed $K=10$. NN trained with $N=1000$ samples, lr $10^{-3}$.
(b) Varying $K$. Each run uses 1000 epochs, $N=1000$, lr $10^{-3}$.
Figure 2: Empirical convergence of ISL’s Bernstein approximation (cf. Eq. 3). The solid blue curve shows the mean theoretical upper bound $\|q_{K}-1\|_{\infty}\leq(K+1)^{2}d_{K}$, and the dashed red curve shows the observed $\|q-1\|_{\infty}$.
Theorem 4.4 (Explicit density approximation).
Let $p,\tilde{p}\in C(\mathbb{R})$ with $\tilde{p}(x)>0$ for all $x\in\mathbb{R}$, and let $\tilde{F}(x)$ be the cdf of $\tilde{p}$. Define
$p_{K}(x):=\tilde{p}(x)\,\sum_{m=0}^{K}\mathbb{Q}_{K}(m)\,\tilde{b}_{m,K}\!\bigl{(}\tilde{F}(x)\bigr{)}.$ (4)
Then for every $x\in\mathbb{R}$,
$\lim_{K\to\infty}p_{K}(x)=p(x).$
Proof.
See Proof in Appendix B. ∎
Remark 4.3.
In practice, one draws latent samples $z_{1},\dots,z_{N}\overset{\mathrm{i.i.d.}}{\sim}p_{z}$ and computes $x_{i}=f(z_{i}),$ where $f$ is the neural network pushing $p_{z}$ forward to $\tilde{p}$. One then forms the empirical cdf and density estimates
$\widehat{\tilde{F}}(x)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{x_{i}\leq x\},\qquad\widehat{\tilde{p}}(x)=\frac{\widehat{\tilde{F}}(x+\delta)-\widehat{\tilde{F}}(x-\delta)}{2\delta}.$
Substituting these into Equation (4) yields the Monte Carlo approximation
$\widehat{p}_{K}(x)=\widehat{\tilde{p}}(x)\sum_{m=0}^{K}\mathbb{Q}_{K}(m)\,\tilde{b}_{m,K}\bigl{(}\widehat{\tilde{F}}(x)\bigr{)}.$ (5)
In Figure 3 we illustrate ISL’s capability—via its Bernstein polynomial approximation—to recover the true density in both one-dimensional and two-dimensional settings. Figure 3(a) compares the ground-truth mixture Gaussian (red) with dual-ISL estimates at $K=2$ (light blue) and $K=15$ (dark blue), while Figure 3(b) overlays the estimated density contours on the two-moons sample scatter. Additional experiments and implementation details, are provided in Appendix D.5.
(a) Comparison of the true $\mathrm{Cauchy}(1,2)$ density (red), the dual-ISL estimate for $K=10$ (blue), and the kernel density estimate (green).
(b) Two-moons target: sample scatter (blue points) with dual-ISL density overlay.
Figure 3: Dual-ISL density estimation results. (a) On a 1D Cauchy target, dual-ISL (blue) closely matches the true density (red) and outperforms the KDE baseline (green). (b) On a 2D two-moons dataset, dual-ISL accurately captures the manifold structure, with learned contours aligning tightly with the sample cloud.
5 Sliced multivariate ISL via Bernstein polynomial approximation
When the data are multidimensional, the target $p(x)$ is a pdf on $\mathbb{R}^{d}$, with $d>1$, and there is no finite set of univariate statistics that uniquely characterizes an arbitrary density (cf. Theorem 2.3 in 1D). Instead, we employ a sliced strategy: we assess a $d$-dimensional distribution by projecting it onto many random directions, computing the one-dimensional ISL discrepancy along each slice, and then averaging these values [de2024robust, kolouri2019generalized].
One–dimensional projected statistic
For any unit vector in the $d$-dimensional sphere $s\in\mathbb{S}^{d}\subset\mathbb{R}^{d+1}$, denote by $s\#p$ the pdf of the one-dimensional projection $y=s^{\top}x$ with associated cdf denoted by $\tilde{F}_{s}$. Then the pmf of order $K$ in direction $s$ is
$\mathbb{Q}_{K}^{s}(n)$ $=\int_{\mathbb{R}}\binom{K}{n}\,[\tilde{F}_{s}(y)]^{n}\,[1-\tilde{F}_{s}(y)]^{K-n}\,(s\#p)(y)\,\mathrm{d}y$
$=\int_{\mathbb{R}}\binom{K}{n}t^{n}(1-t)^{K-n}\dfrac{s\#p(\tilde{F}_{s}^{-1}(t))}{s\#\tilde{p}(\tilde{F}_{s}^{-1}(t))}\mathrm{d}t=\langle b_{n,K},q^{s}\circ\tilde{F}_{s}^{-1}\rangle_{L^{2}(\mathbb{R})},\qquad n=0,\ldots,K.$
where we have denoted by $q^{s}(x)=s\#q(x)=\dfrac{s\#p}{s\#\tilde{p}}(x)$ the push-forward of the quotient $q$ by the linear transformation $s$.
Sliced ISL divergence
We then define the sliced ISL discrepancy by integrating over the unit sphere,
$d_{K}^{\mathbb{S}^{d}}(p,\tilde{p})\;=\;\int_{\mathbb{S}^{d}}d_{K}\bigl{(}s\#p,\;s\#\tilde{p}\bigr{)}\,\mathrm{d}s,$ (6)
where $d_{K}$ is the discrepancy in Definition 2. In practice, to approximate the integral in Equation 6, one randomly samples a finite set of directions $\{s_{\ell}\}_{\ell=1}^{L}$ and averages the resulting evaluations.
The following Theorem is derived using the bounds of Equation 3, and shows that, under mild smoothness assumptions on $q(x)=p(x)/\tilde{p}(x)$, if $\lim_{K\to\infty}d^{\mathbb{S}^{d}}_{K}(p,\tilde{p})=0$ then every one-dimensional projected ratio $q^{s}$ converges uniformly to 1. By the Cramér–Wold theorem [billingsley2017probability, Thm. 29.4], this ensures that $p=\tilde{p}$ on $\mathbb{R}^{d}$, and hence $d_{\mathbb{S}^{d}}^{K}$ becomes a proper divergence as $K\to\infty$.
Theorem 5.1 (Uniform convergence under slicing).
Let $p,\tilde{p}\in C^{2}\left(\mathbb{R}^{d}\right)$. Then there is a constant $C_{d}=\mathcal{L}(\mathbb{S}^{d})$ such that
  1. (i)
[…]
  2. (ii)
    $\sup_{s\in\mathbb{S}^{d}}\|q^{s}-1\|_{\infty}\leq(K+1)^{2}\,\sup_{s\in\mathbb{S}^{d}}d_{K}\bigl{(}s\#p,s\#\tilde{p}\bigr{)}+\frac{\|\nabla^{2}q\|_{\infty}}{8K}.$
Here $\|\nabla^{2}q\|_{\infty}=\sup_{x\in[0,1]^{d}}\|\nabla^{2}q(x)\|$ and $\mathcal{L}(\mathbb{S}^{d})$ is the surface measure of the sphere.
Proof.
See Appendix C. ∎
Since $s\mapsto s\#(\cdot)$ is linear, compactness of $\mathbb{S}^{d}$ plus Theorems 3.13.2 imply that $\;(p,\tilde{p})\mapsto d_{K}^{\mathbb{S}^{d}}(p,\tilde{p})$ is continuous and convex in its first argument. Consequently, by interchanging the roles of the model and target distributions in the slicing framework we obtain a sliced dual-ISL method that retains both convexity and differentiability (almost everywhere) under mild smoothness of the network parameters. Pseudocode for its implementation is given in Appendix E.
6 Summary and concluding remarks
In this paper, we introduced dual-ISL, a novel likelihood-free objective that significantly advances the training of implicit generative models. By interchanging the roles of the target and model distributions within the Invariant Statistical Loss (ISL) framework, dual-ISL provides a convex optimization problem in the space of model densities, addressing common challenges like instability, non-convexity, and mode collapse prevalent in existing methods.
A central theoretical contribution is the interpretation of dual-ISL as an explicit $L^{2}$-projection of the push-forward density ratio $q\circ\tilde{F}^{-1}$ onto a Bernstein polynomial basis. This innovative projection approach yields an explicit closed-form approximation of the density ratio, enabling efficient and analytically tractable density evaluation—a capability traditionally missing in implicit modeling. We derived precise error bounds and convergence rates leveraging classical results from polynomial approximation theory, thus ensuring both theoretical rigor and practical stability.
We further generalized dual-ISL to multivariate distributions through a sliced projection methodology, maintaining convexity, continuity, and analytic tractability in higher-dimensional settings. Empirically, dual-ISL consistently demonstrated improved convergence, smoother training dynamics, and reduced mode collapse compared to classical ISL, GAN variants (including WGAN and MMD-GAN), and normalizing flow baselines across a variety of synthetic benchmarks.
In conclusion, dual-ISL bridges an important gap in implicit generative modeling by providing both strong theoretical foundations and practical advantages. Future directions include exploring adaptive slicing strategies, extending the theoretical analysis to broader classes of polynomial bases, and applying dual-ISL to large-scale generative modeling tasks in diverse real-world domains.
[…]
Appendix A Proofs of Theorems Section 3
In this appendix, we establish three key analytic properties of the rank-based divergence $d_{K}$
  • Continuity in its first argument under weak convergence (Theorem 3.1).
  • Continuity in its second argument under $L^{1}$ norm (Theorem A.1).
  • Convexity in its first argument (Theorem 3.2).
A.1 Proof of Theorem 3.1
Proof of Theorem 3.1.
  1. (i)
    For each fixed $m\in\{0,1,\dots,K\}$, define
    $h_{m}(y)=\binom{K}{m}\,[\tilde{F}(y)]^{m}\,[1-\tilde{F}(y)]^{K-m}.$
[…]
    Taking $h(y)=h_{m}(y)$ yields
[…]
  2. (ii)
    The discrepancy between $p$ and $\tilde{p}$ is
    $d_{K}(p,\tilde{p})=\frac{1}{K+1}\sum_{m=0}^{K}\left|\frac{1}{K+1}-\mathbb{Q}_{K}(m)\right|.$
    Since for each $m$ we have shown that
    $\lim_{n\to\infty}\mathbb{Q}_{K}^{(n)}(m)=\mathbb{Q}_{K}(m),$
    and because the absolute value function is continuous, it follows that
    $\lim_{n\to\infty}\left|\frac{1}{K+1}-\mathbb{Q}_{K}^{(n)}(m)\right|=\left|\frac{1}{K+1}-\mathbb{Q}_{K}(m)\right|.$
    As the sum is finite (from $m=0$ to $K$), we can exchange the limit and the summation to conclude that
    $\lim_{n\to\infty}d_{K}(p_{n},\tilde{p})$ $=\frac{1}{K+1}\sum_{m=0}^{K}\lim_{n\to\infty}\left|\frac{1}{K+1}-\mathbb{Q}_{K}^{(n)}(m)\right|$
    $=\frac{1}{K+1}\sum_{m=0}^{K}\left|\frac{1}{K+1}-\mathbb{Q}_{K}(m)\right|=d_{K}(p,\tilde{p}).$
A.2 Proof of Theorem A.1
We now state and prove that the divergence $d_{K}$ is continuous with respect to its second argument in the $L^{1}$ norm.
Theorem A.1 (Continuity in the second argument).
Let $p$ and $\{\tilde{p}_{n}\}_{n\geq 1}$ be continuous densities on $\mathbb{R}$ such that $\tilde{p}_{n}\to\tilde{p}$ in the $L^{1}$ norm. Then, the discrepancy function $d_{K}$ is continuous in its second argument, i.e.
$\lim_{n\to\infty}d_{K}\bigl{(}p,\tilde{p}_{n}\bigr{)}\;=\;d_{K}(p,\tilde{p}).$
Proof.
Recall that for any pair of densities $p,\tilde{p}$ we define (in Section 4)
$\Phi_{K}(p,\tilde{p})=\bigl{(}\mathbb{Q}_{K}(0),\,\mathbb{Q}_{K}(1),\,\dots,\,\mathbb{Q}_{K}(K)\bigr{)}\;\in\;\mathbb{R}^{K+1},$
and we write the $m$th component of $\Phi_{K}(p,\tilde{p})$ as
$[\Phi_{K}(p,\tilde{p})]_{m}\;=\;\mathbb{Q}_{K}(m).$
Thus, if $\tilde{p}_{n}$ has cdf $\tilde{F}_{n}(y)$, then by definition
$[\Phi_{K}(p,\tilde{p}_{n})]_{m}=\int_{\mathbb{R}}\binom{K}{m}\,\bigl{[}\tilde{F}_{n}(y)\bigr{]}^{m}\bigl{[}1-\tilde{F}_{n}(y)\bigr{]}^{K-m}\,p(y)\,\mathrm{d}y.$
Hence,
$\|\Phi_{K}(p,\tilde{p}_{n})$ $-\Phi_{K}(p,\tilde{p})\|_{\ell^{1}}$
$=$ $\sum_{m=0}^{K}\left|\int_{\mathbb{R}}\binom{K}{m}\,\Bigl{(}[\tilde{F}_{n}(y)]^{m}[1-\tilde{F}_{n}(y)]^{K-m}\;-\;[\tilde{F}(y)]^{m}[1-\tilde{F}(y)]^{K-m}\Bigr{)}p(y)\,\mathrm{d}y\right|$
and by the triangle inequality we have
$\|\Phi_{K}(p,\tilde{p}_{n})$ $-\Phi_{K}(p,\tilde{p})\|_{\ell^{1}}\,\,$
$\leq$ $\sum_{m=0}^{K}\binom{K}{m}\int_{\mathbb{R}}\Bigl{|}[\tilde{F}_{n}(y)]^{m}[1-\tilde{F}_{n}(y)]^{K-m}-[\tilde{F}(y)]^{m}[1-\tilde{F}(y)]^{K-m}\Bigr{|}p(y)\,\mathrm{d}y.$
The function
$f(a)\;=\;a^{m}\,(1-a)^{K-m},$
is Lipschitz on $[0,1]$ with a Lipschitz constant $C_{K,m}<\infty$. As a consequence,
$\bigl{|}f(\tilde{F}_{n}(y))-f(\tilde{F}(y))\bigr{|}\leq C_{K}\,\bigl{|}\tilde{F}_{n}(y)-\tilde{F}(y)\bigr{|},$
where $C_{K}=\sup_{0\leq m\leq K}C_{K,m}<\infty$, and therefore,
$\|\Phi_{K}(p,\tilde{p}_{n})$ $-\Phi_{K}(p,\tilde{p})\|_{\ell^{1}}\,\leq\,C^{\prime}_{K}\int_{\mathbb{R}}\bigl{|}\tilde{F}_{n}(y)-\tilde{F}(y)\bigr{|}\,p(y)\,\mathrm{d}y,$ (7)
where the constant $C^{\prime}_{K}=C_{K}\sum_{m=0}^{K}\binom{K}{m}<\infty$ depends only on $K$.
Note that
$\tilde{F}_{n}(y)\;-\;\tilde{F}(y)\;=\;\int_{-\infty}^{y}\bigl{(}\tilde{p}_{n}(t)-\tilde{p}(t)\bigr{)}\,\mathrm{d}t,$
hence
$\bigl{|}\tilde{F}_{n}(y)-\tilde{F}(y)\bigr{|}\;\leq\;\int_{\mathbb{R}}\bigl{|}\tilde{p}_{n}(t)-\tilde{p}(t)\bigr{|}\,\mathrm{d}t\;=\;\|\tilde{p}_{n}-\tilde{p}\|_{L^{1}},$
and, as a consequence,
$\int_{\mathbb{R}}\bigl{|}\tilde{F}_{n}(y)-\tilde{F}(y)\bigr{|}p(y)\,\mathrm{d}y\;\leq\;\|\tilde{p}_{n}-\tilde{p}\|_{L^{1}}\int_{\mathbb{R}}p(y)\,\mathrm{d}y\;=\;\|\tilde{p}_{n}-\tilde{p}\|_{L^{1}},$ (8)
since $\int_{\mathbb{R}}p(y)\,\mathrm{d}y=1$.
Combining Equation 7 and 8 we arrive at
$\|\Phi_{K}(p,\tilde{p}_{n})$ $-\Phi_{K}(p,\tilde{p})\|_{\ell^{1}}\;\leq\;C^{\prime}_{K}\,\|\tilde{p}_{n}-\tilde{p}\|_{L^{1}}\;\longrightarrow\;0\quad\text{as }n\to\infty.$
Finally, since
$d_{K}\bigl{(}p,\tilde{p}_{n}\bigr{)}\;=\;\frac{1}{K+1}\;\Bigl{\|}\Phi_{K}(p,\tilde{p}_{n})-\mathbb{U}_{K}\Bigr{\|}_{\ell^{1}},$
we have
$\bigl{|}d_{K}\bigl{(}p,\tilde{p}_{n}\bigr{)}\;-\;d_{K}\bigl{(}p,\tilde{p}\bigr{)}\bigr{|}\;=\;\frac{1}{K+1}\bigl{|}\|\Phi_{K}(p,\tilde{p}_{n})-\mathbb{U}_{K}\|_{\ell^{1}}-\|\Phi_{K}(p,\tilde{p})-\mathbb{U}_{K}\|_{\ell^{1}}\bigr{|}.$ (9)
By the triangle inequality for $\ell^{1}$,
$\left|\|\Phi_{K}(p,\tilde{p}_{n})-\mathbb{U}_{K}\|_{\ell^{1}}-\|\Phi_{K}(p,\tilde{p})-\mathbb{U}_{K}\|_{\ell^{1}}\right|\;\leq\;\|\Phi_{K}(p,\tilde{p}_{n})-\Phi_{K}(p,\tilde{p})\|_{\ell^{1}},$ (10)
hence taking 9 and 10 together yields
$\bigl{|}d_{K}\bigl{(}p,\tilde{p}_{n}\bigr{)}\;-\;d_{K}\bigl{(}p,\tilde{p}\bigr{)}\bigr{|}\;\leq\;\frac{1}{K+1}\,\|\Phi_{K}(p,\tilde{p}_{n})-\Phi_{K}(p,\tilde{p})\|_{\ell^{1}}.$
Since $\|\Phi_{K}(p,\tilde{p}_{n})-\Phi_{K}(p,\tilde{p})\|_{\ell^{1}}\to 0$, it follows that
$\lim_{n\to\infty}d_{K}\bigl{(}p,\tilde{p}_{n}\bigr{)}\;=\;d_{K}\bigl{(}p,\tilde{p}\bigr{)},$
A.3 Proof of Theorem 3.2
Finally we give the proof of the convexity of $d_{K}$ w.r.t. the first argument.
Proof of Theorem 3.2.
Let $p_{1},p_{2}$ be two densities and $\lambda\in[0,1]$. Define
$p(y)=\lambda p_{1}(y)+(1-\lambda)p_{2}(y).$
Since
$[\Phi_{K}(p,\tilde{p})]_{n}=\int_{\mathbb{R}}\binom{K}{n}\,[\tilde{F}(y)]^{n}[1-\tilde{F}(y)]^{K-n}\,p(y)\,\mathrm{d}y,$
linearity of the integral gives
$\Phi_{K}(p,\tilde{p})=\lambda\,\Phi_{K}(p_{1},\tilde{p})+(1-\lambda)\,\Phi_{K}(p_{2},\tilde{p}).$
Hence, for each $n$,
$\left|[\Phi_{K}(p,\tilde{p})]_{n}-\tfrac{1}{K+1}\right|$ $=$ $\left|\lambda\bigl{(}[\Phi_{K}(p_{1},\tilde{p})]_{n}-\tfrac{1}{K+1}\bigr{)}+(1-\lambda)\bigl{(}[\Phi_{K}(p_{2},\tilde{p})]_{n}-\tfrac{1}{K+1}\bigr{)}\right|$
$\leq$ $\lambda\,\left|[\Phi_{K}(p_{1},\tilde{p})]_{n}-\tfrac{1}{K+1}\right|+(1-\lambda)\,\left|[\Phi_{K}(p_{2},\tilde{p})]_{n}-\tfrac{1}{K+1}\right|.$
Summing over $n=0,\dots,K$ and dividing by $K+1$ yields
$d_{K}(p,\tilde{p})=\frac{1}{K+1}\sum_{n=0}^{K}\bigl{|}\Phi_{K}(p,\tilde{p})_{n}-\tfrac{1}{K+1}\bigr{|}\leq\lambda\,d_{K}(p_{1},\tilde{p})+(1-\lambda)\,d_{K}(p_{2},\tilde{p}).$
Appendix B Proofs of Section 4
In this appendix we give complete proofs for the theorems and claims in Sections 4.
  • Theorem B.1. Characterizes the binomial mapping $\Phi_{K}$, showing it is well-defined, linear in its first argument, bounded, and continuous under mild regularity conditions.
  • Theorem 4.1. Establishes the Riesz representation of $\Phi_{K}$, expressing each probability mass $\mathbb{Q}_{K}(n)$ as an $L^{2}$ inner product with a Bernstein basis function.
  • Theorem 4.3. Provides a uniform bound on the deviation $\|q_{K}-1\|_{\infty}$ in terms of the discrepancy $d_{K}(p,\tilde{p})$.
  • Theorem 4.4. Derives an explicit Bernstein-based representation for a push-forward density $p_{K}$ via a continuously differentiable map.
B.1 Properties of the map $\Phi_{K}$
Theorem B.1.
Let $p$ and $\tilde{p}$ be pdfs on $\mathbb{R}$, with cdfs $F$ and $\tilde{F}$, respectively. For each integer $K\geq 1$, recall that
$\Phi_{K}(p,\tilde{p})=\bigl{(}\mathbb{Q}_{K}(0),\,\mathbb{Q}_{K}(1),\,\dots,\,\mathbb{Q}_{K}(K)\bigr{)}\;\in\;\mathbb{R}^{K+1},$
where
$\mathbb{Q}_{K}(n)\;=\;\int_{\mathbb{R}}\binom{K}{n}\,\bigl{[}\tilde{F}(y)\bigr{]}^{n}\bigl{[}1-\tilde{F}(y)\bigr{]}^{K-n}\,p(y)\,\mathrm{d}y,\quad n=0,\ldots,K.$ (11)
Then the following properties hold
  1. (i)
    Well-definedness. For each fixed pair $(p,\tilde{p})$ and integer $K$, the integral on the right -hand side of Equation 11 uniquely determines a pmf $\mathbb{Q}_{K}$.
  2. (ii)
    Non-surjectivity. Let $\Delta^{K}$ be the set of all pmfs on $\{0,1,\ldots,K\}$. The image of $\Phi_{K}$ is strictly contained in $\Delta^{K}$.
  3. (iii)
[…]
  4. (iv)
    Linearity. The operator $\Phi_{K}$ is linear in its first argument. Thus, for any $\alpha\in[0,1]$,
    $\Phi_{K}\big{(}\alpha p_{1}+(1-\alpha)p_{2},\tilde{p}\big{)}=\alpha\,\Phi_{K}(p_{1},\tilde{p})+(1-\alpha)\,\Phi_{K}(p_{2},\tilde{p}).$
  5. (v)
    Bounded operator. $\Phi_{K}$ is a bounded operator from the space of continuous pdfs $p$ on $\mathbb{R}$ to the space of pmfs $\mathbb{Q}_{K}$ on $\{0,1,\ldots,K\}$. Specifically
    $\left\lVert\!\!\left\lVert\!\!\left\lVert\Phi_{K}\right\rVert\!\!\right\rVert\!\!\right\rVert=\sup_{\|p\|_{L^{1}}\leq 1}\|\Phi_{K}(p,\tilde{p})\|_{\text{TV}}=\|\mathbb{Q}_{K}\|_{TV}=1,$
[…]
Proof of Theorem B.1.
  1. (i)
    Fix $K\geq 0$ and recall that
    $\mathbb{Q}_{K}(n)=\int_{\mathbb{R}}\binom{K}{n}\,[\tilde{F}(y)]^{n}\,[1-\tilde{F}(y)]^{K-n}\,p(y)\,\mathrm{d}y,\qquad n=0,\dots,K.$
    Non-negativity. Since $p(y)\geq 0$ and each Bernstein-integrand $\binom{K}{n}[\tilde{F}(y)]^{n}[1-\tilde{F}(y)]^{K-n}\geq 0$, it follows that $\mathbb{Q}_{K}(n)\geq 0$ for every $n$.
    Normalization. By Fubini’s theorem we may interchange sum and integral, i.e.,
    $\sum_{n=0}^{K}\mathbb{Q}_{K}(n)=\int_{\mathbb{R}}p(y)\sum_{n=0}^{K}\binom{K}{n}[\tilde{F}(y)]^{n}[1-\tilde{F}(y)]^{K-n}\,\mathrm{d}y.$
    But the inner sum is $(\tilde{F}(y)+(1-\tilde{F}(y)))^{K}=1^{K}=1$ by the binomial theorem, hence
    $\sum_{n=0}^{K}\mathbb{Q}_{K}(n)=\int_{\mathbb{R}}p(y)\,\mathrm{d}y=1.$
    Uniqueness. Each $\mathbb{Q}_{K}(n)$ is defined by a single integral depending only on $p$ and $\tilde{F}$. Thus the mapping $\Phi_{K}:(p,\tilde{p})\mapsto(\mathbb{Q}_{K}(0),\ldots,\mathbb{Q}_{K}(K))$ is well-defined and unique for each choice of $(p,\tilde{p})$ and $K$.
  2. (ii)
    To see that $\Phi_{K}$ is not surjective, it suffices to exhibit a pmf in $\Delta^{K}$ that cannot arise from any $(p,\tilde{p})$. We do this for $K=2$.
    Define
    $\mathbb{Q}=(0,1,0)\in\Delta^{2},$
    so that $\mathbb{Q}(0)=0$, $\mathbb{Q}(1)=1$, and $\mathbb{Q}(2)=0$. Suppose, for the sake of contradiction, that there exist densities $p,\tilde{p}$ on $\mathbb{R}$ with cdf $\tilde{F}$ such that $\Phi_{2}(p,\tilde{p})=\mathbb{Q}$. Then by definition
    $\mathbb{Q}(0)$ $=\int_{\mathbb{R}}[1-\tilde{F}(y)]^{2}\,p(y)\,\mathrm{d}y=0,$
    $\mathbb{Q}(1)$ $=2\int_{\mathbb{R}}\tilde{F}(y)\,[1-\tilde{F}(y)]\,p(y)\,\mathrm{d}y=1,$
    $\mathbb{Q}(2)$ $=\int_{\mathbb{R}}[\tilde{F}(y)]^{2}\,p(y)\,\mathrm{d}y=0.$
    The first and third equations force
[…]
    hence $\tilde{F}(y)=1$ and $\tilde{F}(y)=0$ $p$-a.e. Since $p$ is a probability density, its support has positive measure, so we cannot have $\tilde{F}\equiv 1$ and $\tilde{F}\equiv 0$ on that support. This contradiction shows no $(p,\tilde{p})$ can produce $\mathbb{Q}=(0,1,0)$. Therefore $\Phi_{2}$, and hence $\Phi_{K}$ for general $K$, fails to be surjective.
  3. (iii)
    Let $p_{i}\to p$ in $L^{1}(\mathbb{R})$ and let $\tilde{p}_{j}$ have cdfs $\tilde{F}_{j}\to\tilde{F}$ uniformly on $\mathbb{R}$. We show
[…]
    By the triangle inequality,
    $\bigl{\|}\Phi_{K}(p_{i},\tilde{p}_{n})-\Phi_{K}(p,\tilde{p})\bigr{\|}_{\ell^{1}}$ $\leq\bigl{\|}\Phi_{K}(p_{i},\tilde{p}_{n})-\Phi_{K}(p_{i},\tilde{p})\bigr{\|}_{\ell^{1}}$
    $\quad+\;\bigl{\|}\Phi_{K}(p_{i},\tilde{p})-\Phi_{K}(p,\tilde{p})\bigr{\|}_{\ell^{1}},$ (12)
    and we handle each term separately.
    Continuity in $\tilde{p}$. Fix $p_{i}$. For each $m=0,\dots,K$, set
    $g_{j}(y)=[\tilde{F}_{j}(y)]^{m}\,[1-\tilde{F}_{j}(y)]^{K-m},\quad g(y)=[\tilde{F}(y)]^{m}\,[1-\tilde{F}(y)]^{K-m}.$
    Uniform convergence $\|\tilde{F}_{j}-\tilde{F}\|_{\infty}\to 0$ implies $\|g_{j}-g\|_{\infty}\to 0$. Hence for each coordinate $m$ we have
    $\bigl{|}\Phi_{K}(p_{i},\tilde{p}_{n})_{m}-\Phi_{K}(p_{i},\tilde{p})_{m}\bigr{|}\leq\binom{K}{m}\int_{\mathbb{R}}|g_{i}(y)-g(y)|\,p_{i}(y)\,\mathrm{d}y\leq\binom{K}{m}\|g_{n}-g\|_{\infty},$
    and summing over $m$ yields
[…]
    Continuity in $p$. Fix $\tilde{p}$ (and write $\tilde{F}$ for its cdf). For each $m=0,\dots,K$,
    $\bigl{[}\Phi_{K}(p_{i},\tilde{p})\bigr{]}_{m}-\bigl{[}\Phi_{K}(p,\tilde{p})\bigr{]}_{m}=\binom{K}{m}\int_{\mathbb{R}}[\tilde{F}(y)]^{m}[1-\tilde{F}(y)]^{K-m}\bigl{(}p_{i}(y)-p(y)\bigr{)}\,\mathrm{d}y.$
    Taking absolute values and using $\int|p_{i}-p|=\|p_{i}-p\|_{L^{1}}$ yields
    $\bigl{|}\Phi_{K}(p_{i},\tilde{p})_{m}-\Phi_{K}(p,\tilde{p})_{m}\bigr{|}\;\leq\;\binom{K}{m}\,\|p_{i}-p\|_{L^{1}}.$
    Summing over $m$ gives
    $\bigl{\|}\Phi_{K}(p_{i},\tilde{p})-\Phi_{K}(p,\tilde{p})\bigr{\|}_{\ell^{1}}=$ $\sum_{m=0}^{K}\bigl{|}\Phi_{K}(p_{i},\tilde{p})_{m}-\Phi_{K}(p,\tilde{p})_{m}\bigr{|}$
[…]
    Since $p_{i}\to p$ in $L^{1}$, the right-hand side tends to zero. Hence $\Phi_{K}(\cdot,\tilde{p})$ is continuous in its first argument.
    Combining in Equations (iii), 13 and (iii) yields
[…]
    i.e. $\Phi_{K}$ is jointly continuous in $(p,\tilde{p})$.
  4. (iv)
    Linearity in $p$. Let $p_{1},p_{2}$ be two probability densities on $\mathbb{R}$ and $\alpha\in[0,1]$. Set
    $p=\alpha\,p_{1}+(1-\alpha)\,p_{2}.$
    Then for each $n=0,\dots,K$,
    $\Phi_{K}(p,\tilde{p})_{n}=\binom{K}{n}\int_{\mathbb{R}}[\tilde{F}(y)]^{n}\,[1-\tilde{F}(y)]^{K-n}\,p(y)\,\mathrm{d}y.$
    By linearity of the integral,
    $\Phi_{K}(p,\tilde{p})_{n}$ $=\binom{K}{n}\!\int_{\mathbb{R}}[\tilde{F}]^{n}[1-\tilde{F}]^{K-n}\bigl{(}\alpha p_{1}+(1-\alpha)p_{2}\bigr{)}\,\mathrm{d}y$
    $=\alpha\,\Phi_{K}(p_{1},\tilde{p})_{n}+(1-\alpha)\,\Phi_{K}(p_{2},\tilde{p})_{n}.$
    Since this holds for every coordinate $n$, we conclude
    $\Phi_{K}\bigl{(}\alpha p_{1}+(1-\alpha)p_{2},\;\tilde{p}\bigr{)}=\alpha\,\Phi_{K}(p_{1},\tilde{p})+(1-\alpha)\,\Phi_{K}(p_{2},\tilde{p}),$
    i.e. $\Phi_{K}(\cdot,\tilde{p})$ is linear.
  5. (v)
    Boundedness and operator norm. Recall the total-variation norm on $\mathbb{R}^{K+1}$ is just the $\ell^{1}$-norm. For any pair of densities $p,\tilde{p}$,
    $\|\Phi_{K}(p,\tilde{p})\|_{\mathrm{TV}}=\sum_{n=0}^{K}[\Phi_{K}(p,\tilde{p})]_{n}=\int_{\mathbb{R}}p(y)\sum_{n=0}^{K}\binom{K}{n}[\tilde{F}(y)]^{n}[1-\tilde{F}(y)]^{K-n}\,\mathrm{d}y.$
    By the binomial theorem the inner sum is $\bigl{(}\tilde{F}+(1-\tilde{F})\bigr{)}^{K}=1$. Hence
    $\|\Phi_{K}(p,\tilde{p})\|_{\mathrm{TV}}=\int_{\mathbb{R}}p(y)\,\mathrm{d}y=\|p\|_{L^{1}}.$
    Taking the supremum over all $p$ with $\|p\|_{L^{1}}\leq 1$ shows
    $\left\lVert\!\!\left\lVert\!\!\left\lVert\Phi_{K}\right\rVert\!\!\right\rVert\!\!\right\rVert:=\sup_{\|p\|_{L^{1}}\leq 1}\|\Phi_{K}(p,\tilde{p})\|_{\mathrm{TV}}=\sum_{n=0}^{K}\mathbb{Q}_{K}(n)=1.$
    Thus $\Phi_{K}$ is a bounded linear operator with $\|\Phi_{K}\|=1$.
B.2 Proof of Theorem 4.1
Proof of Theorem 4.1.
Recall that, by Theorem B.1, for each fixed $\tilde{p}$, the map
$\Phi_{K}(\cdot,\tilde{p}):C(\mathbb{R})\;\longrightarrow\;\mathbb{R}^{K+1}$
is linear. To exhibit its Riesz representation, it suffices to find, for each $n=0,1,\dots,K$, a function $f_{n}(y)\in L^{2}(\mathbb{R})$ such that
$[\Phi_{K}(p,\tilde{p})]_{n}\;=\;\mathbb{Q}_{K}(n)\;=\;\int_{\mathbb{R}}f_{n}(y)\,p(y)\,\mathrm{d}y\;=\;\bigl{\langle}f_{n},\,p\bigr{\rangle}_{L^{2}(\mathbb{R})}.$
However, by definition of $\Phi_{K}$,
$\mathbb{Q}_{K}(n)\;=\;\binom{K}{n}\int_{\mathbb{R}}\bigl{[}\tilde{F}(y)\bigr{]}^{n}\,\bigl{[}1-\tilde{F}(y)\bigr{]}^{K-n}\,p(y)\,\mathrm{d}y=\int_{\mathbb{R}}\left(b_{n,K}\circ\tilde{F}\right)(y)\,p(y)\,\mathrm{d}y,$
which is exactly $\langle\,b_{n,K}\circ\tilde{F},\;p\rangle$ for the Bernstein polynomial of degree $K$. Therefore
$\Phi_{K}(p,\tilde{p})\;=\;\bigl{(}\langle f_{0},p\rangle,\,\langle f_{1},p\rangle,\,\dots,\,\langle f_{K},p\rangle\bigr{)},$
and the Riesz theorem yields the claimed representation, with $\{f_{n}\}_{n=0}^{K}$ playing the role of the dual elements.
Now assume $\tilde{p}(x)>0$ for all $x$. Then $\tilde{F}^{\prime}(x)=\tilde{p}(x)>0$. So $\tilde{F}$ is continuous and strictly increasing. Hence $\tilde{F}$ is a bijection and admits the inverse $\tilde{F}^{-1}$.
Setting $q(x)=p(x)/\tilde{p}(x)$ and changing variables $t=\tilde{F}(x)$ (so $x=\tilde{F}^{-1}(t)$ and $\mathrm{d}t=\tilde{p}(\tilde{F}^{-1}(t))\,\mathrm{d}x$) yields
$\mathbb{Q}_{K}(n)=\int_{\mathbb{R}}b_{n,K}\bigl{(}\tilde{F}(x)\bigr{)}\,p(x)\,\mathrm{d}x=\int_{0}^{1}b_{n,K}(t)\,q\bigl{(}\tilde{F}^{-1}(t)\bigr{)}\,\mathrm{d}t=\bigl{\langle}b_{n,K},\,q\circ\tilde{F}^{-1}\bigr{\rangle}_{L^{2}([0,1])}.$
This completes the proof. ∎
B.3 Proof of Theorem 4.2
Proof of Theorem 4.2.
Existence of the Bernstein expansion. Since $p,\tilde{p}\in C(\mathbb{R})$ and $\tilde{p}(x)>0$ for all $x$, the ratio $q(x)\;=\;\frac{p(x)}{\tilde{p}(x)}$ is well-defined and continuous on $\mathbb{R}$. Moreover, the cdf $\tilde{F}$ is $C^{1}$ with $\tilde{F}^{\prime}(x)=\tilde{p}(x)>0$; hence $\tilde{F}$ is strictly increasing and continuous. It follows that $\tilde{F}^{-1}$ exists and is continuous . Therefore, setting
$\tilde{q}(t)\;=\;q\bigl{(}\tilde{F}^{-1}(t)\bigr{)},$
we obtain $\tilde{q}\in C([0,1])$.
Since $\{b_{n,K}\}_{n\geq 0}$ is a Schauder basis of $C([0,1])$ (see [lorentz2012bernstein, Theorem 1.1.1] and [megginson2012introduction, Chapter 4]), there are unique coefficients $\{\alpha_{n}\}_{n\geq 0}$ such that
$\tilde{q}(t)=\sum_{n=0}^{\infty}\alpha_{n}\,b_{n,n}(t),\qquad t\in[0,1].$
Truncating at degree $K$ yields
$\tilde{q}_{K}(t):=\sum_{n=0}^{K}\alpha_{n}\,b_{n,K}(t).$
Dual-basis (projection) representation. Let
$V_{K}=\mathrm{span}\{b_{0,K},\dots,b_{K,K}\}\subset L^{2}([0,1]),$
and define the Gram matrix $G$ with entries
$G_{n,m}=\bigl{\langle}b_{n,K},\,b_{m,K}\bigr{\rangle}=\int_{0}^{1}b_{n,K}(t)\,b_{m,K}(t)\,\mathrm{d}t,\qquad n,m\in\{0,\ldots,K\},$
which is nonsingular (positive-definite). The dual (biorthogonal) basis $\{\tilde{b}_{n,K}\}\subset V_{K}$ is given by (see [jiittler1998dual, Section 2])
$\tilde{b}_{n,K}(t)=\sum_{m=0}^{K}(G^{-1})_{n,m}\,b_{m,K}(t),$
so that $\langle\tilde{b}_{n,K},b_{m,K}\rangle=\delta_{n,m}$.
Let $P_{K}:L^{2}([0,1])\to V_{K}$ the orthogonal projection constructed as $P_{K}(f)\in V_{K}$ that satisfies $f-P_{K}(f)\perp V_{K}$. We write
$P_{K}(f)=\sum_{n=0}^{K}c_{n}\,\tilde{b}_{n,K},$
and impose for each $m$,
$0=\bigl{\langle}f-P_{K}(f),\,b_{m,K}\bigr{\rangle}=\bigl{\langle}f,\,b_{m,K}\bigr{\rangle}-\sum_{n=0}^{K}c_{n}\,\bigl{\langle}\tilde{b}_{n,K},\,b_{m,K}\bigr{\rangle}=\bigl{\langle}f,\,b_{m,K}\bigr{\rangle}-c_{m},$
where $c_{m}=\langle f,b_{m,K}\rangle$. Thus
$P_{K}(f)=\sum_{n=0}^{K}\bigl{\langle}f,b_{n,K}\bigr{\rangle}_{L^{2}([0,1])}\,\tilde{b}_{n,K}.$
Applying this to $f=\tilde{q}$ gives
$\tilde{q}_{K}=P_{K}(\tilde{q})=\sum_{n=0}^{K}\bigl{\langle}\tilde{q},\,b_{n,K}\bigr{\rangle}_{L^{2}([0,1])}\,\tilde{b}_{n,K}(t).$
Identification of the coefficients. Since $\tilde{q}(t)=q(\tilde{F}^{-1}(t))$ and $q(x)=p(x)/\tilde{p}(x)$, the change of variables $t=\tilde{F}(x)$ yields
$\bigl{\langle}\tilde{q},\,b_{n,K}\bigr{\rangle}=\int_{0}^{1}b_{n,K}(t)\,\tilde{q}(t)\,\mathrm{d}t$ $=\int_{\mathbb{R}}b_{n,K}\bigl{(}\tilde{F}(x)\bigr{)}\,q(x)\,\tilde{p}(x)\,\mathrm{d}x$
$=\int_{\mathbb{R}}b_{n,K}\bigl{(}\tilde{F}(x)\bigr{)}\,p(x)\,\mathrm{d}x=\mathbb{Q}_{K}(n).$
Hence
$\tilde{q}_{K}(t)=\sum_{n=0}^{K}\mathbb{Q}_{K}(n)\,\tilde{b}_{n,K}(t).$
B.4 Proof Theorem 4.3
In this section we derive a uniform bound on the Bernstein–truncated ratio $q_{K}$. The main tool is the dual-basis expansion of $\tilde{q}_{K}$ and the fact (Lemma B.1) that the dual Bernstein functions sum to $K+1$.
Lemma B.1.
The dual Bernstein basis functions $\tilde{b}_{m,K}(x)$ satisfy
$\sum_{m=0}^{K}\tilde{b}_{m,K}(x)=K+1.$
Proof.
We begin by showing that the constant vector
$e=(1,1,\dots,1)^{\top}$
is an eigenvector of the Gram matrix $G$ with eigenvalue $\lambda=1/(K+1)$. Writing out the $n$th component of $G\,e$ yields
$(G\,e)_{n}=\sum_{m=0}^{K}G_{nm}\,e_{m}=\sum_{m=0}^{K}\int_{0}^{1}b_{n,K}(x)\,b_{m,K}(x)\,\mathrm{d}x=\int_{0}^{1}b_{n,K}(x)\Bigl{(}\sum_{m=0}^{K}b_{m,K}(x)\Bigr{)}\mathrm{d}x.$
Since the Bernstein polynomials form a partition of unity, $\sum_{m=0}^{K}b_{m,K}(x)=1$ (by the Binomial theorem), and each integrates to $1/(K+1)$ (indeed, $\int_{0}^{1}b_{n,K}(x)\,\mathrm{d}x=\binom{K}{n}B(n+1,K-n+1)=\frac{1}{K+1}\!$ by the Beta-function identity), we obtain
$(G\,e)_{n}=\int_{0}^{1}b_{n,K}(x)\,\mathrm{d}x=\frac{1}{K+1}.$
Hence
$G\,e=\frac{1}{K+1}\,e,$ (15)
i.e. $e$ is an eigenvector with eigenvalue $1/(K+1)$.
Left-multiplying by $G^{-1}$ (positive-definite matrix) on both sides of Equation 15 yields
$G^{-1}e=(K+1)\,e.$
Which implies that
$(G^{-1}e)_{n}=\sum_{m=0}^{K}(G^{-1})_{n,m}=K+1\quad\text{for each }n.$
Finally, by the definition of the dual basis,
$\sum_{m=0}^{K}\tilde{b}_{m,K}(x)$ $=$ $\sum_{m=0}^{K}\sum_{n=0}^{K}(G^{-1})_{n,m}\,b_{n,K}(x)$
$=$ $\sum_{n=0}^{K}b_{n,K}(x)\,\Bigl{(}\sum_{m=0}^{K}(G^{-1})_{nm}\Bigr{)}$
$=$ $(K+1)\sum_{n=0}^{K}b_{n,K}(x)$
$=$ $K+1.$
where we again are used the partition-of-unity property $\sum_{n=0}^{K}b_{n,K}(x)=1$. ∎
Proof of Theorem 4.3.
We have
$\tilde{q}_{K}(t)=\sum_{m=0}^{K}\mathbb{Q}_{K}(m)\tilde{b}_{m,K}(x)=\sum_{m=0}^{K}\left[\frac{1}{K+1}-\left(\frac{1}{K+1}-\mathbb{Q}_{K}(m)\right)\right]\tilde{b}_{m,K}(t).$
Since $\sum_{m=0}^{K}\tilde{b}_{m,K}(x)=K+1$ (see Lemma B.1), we rewrite this explicitly as
$\tilde{q}_{K}(t)=1-\sum_{m=0}^{K}\left(\frac{1}{K+1}-\mathbb{Q}_{K}(m)\right)\tilde{b}_{m,K}(t).$
By the triangle inequality, we have
$|\tilde{q}_{K}(t)-1|\leq\sum_{m=0}^{K}\left|\frac{1}{K+1}-\mathbb{Q}_{K}(m)\right|\;|\tilde{b}_{m,K}(x)|\leq(K+1)\sum_{m=0}^{K}\left|\frac{1}{K+1}-\mathbb{Q}_{K}(m)\right|,$
where the inequality $|\tilde{b}_{m,K}(x)|\leq K+1$ follows from Lemma B.1. Recognizing the definition of $d_{K}(p,\tilde{p})$, we have
$|\tilde{q}_{K}(t)-1|\leq(K+1)^{2}d_{K}(p,\tilde{p}).$
Since $q_{K}(x)=\tilde{q}_{K}\bigl{(}\tilde{F}(x)\bigr{)}$, the supremum over $x\in\mathbb{R}$ coincides with the supremum over $t\in[0,1]$. Hence
$\|q_{K}(x)-1\|_{\infty}\leq(K+1)^{2}d_{K}(p,\tilde{p}).$
B.5 Proof of Theorem 4.4
Proof of Theorem 4.4.
Recall from Theorem 4.2 that
$\tilde{q}_{K}(t)=\sum_{m=0}^{K}\mathbb{Q}_{K}(m)\,\tilde{b}_{m,K}(t)$ (16)
is precisely the degree-$K$ truncation of the Bernstein expansion of $\tilde{q}$. Since $\{b_{n,K}\}_{n\geq 0}$ is a Schauder basis of $\bigl{(}C([0,1]),\|\cdot\|_{\infty}\bigr{)}$, these truncations satisfy
$\|\tilde{q}_{K}-\tilde{q}\|_{\infty}=\sup_{t\in[0,1]}|\tilde{q}_{K}(t)-\tilde{q}(t)|\;\longrightarrow\;0\qquad\text{when }\;K\to\infty,$
i.e. $\tilde{q}_{K}\to\tilde{q}$ uniformly on $[0,1]$ (see [lorentz2012bernstein, Section 1.1]). From Equation. (4) and 16,
$p_{K}(x)=\tilde{p}(x)\,\tilde{q}_{K}\bigl{(}\tilde{F}(x)\bigr{)}.$
Since $\tilde{p}$ and $\tilde{F}$ are continuous, for each fixed $x$ the argument $t_{x}=\tilde{F}(x)\in[0,1]$ is constant, and uniform convergence of $\tilde{q}_{K}$ gives
$\lim_{K\to\infty}\tilde{q}_{K}\bigl{(}t_{x}\bigr{)}=\tilde{q}\bigl{(}t_{x}\bigr{)}.$
Hence
$\lim_{K\to\infty}p_{K}(x)=\tilde{p}(x)\,\lim_{K\to\infty}\tilde{q}_{K}(\tilde{F}(x))=\tilde{p}(x)\,\tilde{q}\bigl{(}\tilde{F}(x)\bigr{)}=\tilde{p}(x)\,\frac{p(x)}{\tilde{p}(x)}=p(x).$
Appendix C Proofs of Section 5
In this subsection we prove Theorem 5.1, which shows that for each projection direction $s$, the Bernstein-projected density ratio $q^{s}$ converges uniformly to 1 at the rate controlled by $d_{K}(p,\tilde{p})$ and the Hessian of $q$.
Proof of Theorem 5.1.
Integrating over $\mathbb{S}^{d}$ on both the left-hand and right hand sides of expression (3) yields
[…]
Since $\tilde{p}(x)>0$ for any $x\in\mathbb{R}$ then $q\in C^{2}(\mathbb{R}^{d})$. Note that for any unit vector $v$, the second directional derivative is $D^{2}_{v,v}q(x)=v^{\top}\nabla^{2}q(x)v$, and hence $|D^{2}_{v,v}q(x)|\leq\|\nabla^{2}q(x)\|$. Taking the supremum over $x$ yields the uniform bound $\|\nabla^{2}q\|_{\infty}$. Therefore, we can further write
$\int_{\mathbb{S}^{d}}\|q^{s}-1\|_{\infty}\,\mathrm{d}s\leq(K+1)^{2}\,d_{K}^{\mathbb{S}^{d}}(p,\tilde{p})+\frac{\|\nabla^{2}q\|_{\infty}}{8K}\,\mathcal{L}(\mathbb{S}^{d}),$
where $\mathcal{L}(\mathbb{S}^{d})$ denotes the Lebesgue measure of $\mathbb{S}^{d}$ and $\|\nabla^{2}q\|_{\infty}$ denotes the supremum over $[0,1]^{d}$ of the operator norm of the Hessian of $q$. Similarly, we obtain
$\sup_{s\in\mathbb{S}^{d}}\left\|q^{s}-1\right\|_{\infty}\leq(K+1)^{2}\,\sup_{s\in\mathbb{S}^{d}}d_{K}(s\#p,s\#\tilde{p})+\frac{\|\nabla^{2}q\|_{\infty}}{8K}.$
Appendix D Supplementary experiments
D.1 Evaluating dual-ISL on 1D Target Distributions
Following the evaluation setup of [zaheer2017gan, de2024training], we draw $N=1000$ i.i.d. samples from each of six one-dimensional benchmark targets. The first three targets are classical pdf (e.g. $\mathcal{N}(0,1)$, Cauchy, and Pareto), while the remaining three are equally-weighted mixtures:
  • Model1: $\tfrac{1}{2}\mathcal{N}(5,2)+\tfrac{1}{2}\mathcal{N}(-1,1)$.
  • Model2: $\tfrac{1}{3}\mathcal{N}(5,2)+\tfrac{1}{3}\mathcal{N}(-1,1)+\tfrac{1}{3}\mathcal{N}(-10,3)$.
  • Model3: $\tfrac{1}{2}\mathcal{N}(-5,2)+\tfrac{1}{2}\mathrm{Pareto}(5,1)$.
All non-diffusion methods (Dual-ISL, ISL, GAN, WGAN and MMD-GAN) use the same generator architecture: a four-layer MLP with ELU activations and layer widths [7, 13, 7, 1]71371[7,\,13,\,7,\,1][ 7 , 13 , 7 , 1 ]. Each is trained for $10^{4}$ epochs using Adam with a fixed learning rate of $10^{-2}$. By contrast, the DDPM baseline employs a four-layer ELU-MLP score network with identical widths, augmented by a 16-dimensional sinusoidal time embedding. It is also trained for $10^{4}$ epochs with Adam (lr = $10^{-2}$) across $T=200$ diffusion steps, where the noise schedule $\{\beta_{t}\}$ is linearly spaced from $10^{-4}$ to $2\times 10^{-2}$. Table 1 summarizes the quantitative results, and the corresponding visualizations are shown in Figure 4.
dual-ISL ISL WGAN MMD-GAN Diffusion
$\mathcal{N}(4,2)$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\text{Cauchy}(1,2)$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\text{Pareto}(1,1)$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\text{Model}_{1}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\text{Model}_{2}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\text{Model}_{3}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: One-dimensional density estimation across six benchmark targets. Each row corresponds to a different true distribution (top to bottom: $\mathcal{N}(4,2)$, Cauchy$(1,2)$, Pareto(1,1)11(1,1)( 1 , 1 ), Model1, Model2, Model3). In each subplot, the red curve shows the ground-truth density and the blue curve shows the model’s estimated density. Columns (left to right) compare dual-ISL, classical ISL, WGAN, MMD-GAN, and a DDPM diffusion baseline, respectively. Dual-ISL more accurately captures multi-modal and heavy-tailed shapes, with reduced mode-collapse and smoother estimates.
Referring to [zaheer2017gan, Theorem 1], in the one-dimensional case any generator that perfectly pushes forward a simple base distribution $p_{z}$ (e.g. uniform or Gaussian) to a target distribution $p$ must implement one of at most two continuous maps. Concretely, if $F_{z}$ and $F$ are the cdfs of $p_{z}$ and $p$, then the two solutions are
$f_{+}(z)\;=\;F^{-1}\bigl{(}F_{z}(z)\bigr{)}\qquad\text{and}\qquad f_{-}(z)\;=\;F^{-1}\bigl{(}1-F_{z}(z)\bigr{)}.$
In practice, a learning algorithm that truly captures the underlying structure of a multimodal or heavy-tailed distribution should recover one of these two “probability-integral-transform” maps.
In Figure 5, we therefore plot, for each method, the learned generator $f_{\theta}(z)$ against the theoretical target map $f_{+}(z)$ for the challenging mixture $\tfrac{1}{2}\mathcal{N}(-5,2)+\tfrac{1}{2}\mathrm{Pareto}(5,1)$ (“Model3”). A perfect fit would lie exactly on the diagonal. As seen, dual-ISL (Subfigure 5(a)) closely tracks the true transformation across the full support of $z$, whereas classical ISL (Subfigure 5(b)) suffers from local distortions around the Pareto tail, and MMD-GAN (Subfigure 5(c)) exhibits even larger deviations—especially where the two modes meet. This visualization makes explicit how dual-ISL more faithfully learns the correct mapping, rather than merely matching moments or densities.
(a) dual-ISL
(b) ISL
(c) MMD-GAN
Figure 5: Comparison of learned generator mappings $f_{\theta}(z)$ against the true probability-integral-transform $f_{+}(z)$ or $f_{-}(z)$ for Model3. Dual-ISL closely follows the ideal map even in the heavy-tailed region, while ISL and MMD-GAN display growing errors, particularly near the mode boundaries and in the Pareto tail.
D.2 Computational Benchmarking of Dual-ISL vs (Classical) ISL
Next, we measure execution time using dedicated benchmarking tools [chen2016robust]. These tools first warm up and calibrate the code to determine the optimal number of iterations per measurement, then execute the code in bundled loops to collect multiple independent samples. We compute statistics—including the minimum, median, mean, and standard deviation—while tracking garbage collection time separately. Table 2 summarizes the results for various target distributions at a fixed $K$. Figure 6 combines two perspectives: (a) total runtime as a function of $K$, and (b) estimation accuracy versus runtime at a fixed $K$. These findings show that Dual-ISL not only runs faster than classical ISL, but also achieves a superior accuracy–runtime balance—and this advantage grows even larger as $K$ increases.
Target ISL Dual-ISL
Median Mean $\pm$ $\sigma$ Memory Median Mean $\pm$ $\sigma$ Memory
$\mathcal{N}(4,2)$ $239.374$ $241.280$ p m 0.434 $17.42$ GiB $22.281$ $22.502$ p m 0.118 $9.27$ GiB
$\mathrm{Pareto}(1,1)$ $238.043$ $239.007$ p m 1.607 $17.42$ GiB $22.207$ $22.109$ p m 0.518 $9.27$ GiB
$\mathrm{Model}_{1}$ $241.683$ $241.905$ p m 1.722 $17.42$ GiB $21.685$ $21.709$ p m 0.044 $9.27$ GiB
$\mathrm{Model}_{2}$ $237.825$ $239.805$ p m 1.832 $17.42$ GiB $22.921$ $22.886$ p m 0.453 $9.27$ GiB
Table 2: Runtime (in seconds) and memory benchmarks for ISL vs. Dual-ISL at fixed $K=10$. Each cell reports median runtime, mean $\pm$ standard deviation, and peak memory usage.
(a) Target density Model2
(b) Target density $\text{Model}_{2}$
Figure 6: Computation time and accuracy trade-off for classical ISL versus Dual-ISL as $K$ increases. (a) Total runtime (in seconds) for 1000 training epochs with batch size $N=1000$ on the $\mathrm{Model}_{2}$ target. (b) Runtime versus mean KSD for both methods with $K=10$ and batch size $N=1000$, illustrating how Dual-ISL maintains lower runtimes improving also accuracy. Dual-ISL consistently outperforms classical ISL in speed, with the gap widening at larger $K$.
(a) $\mathcal{N}(4,2)$
(b) $\text{Model}_{3}$
Figure 7: Training curves for dual-ISL versus classical ISL (dashed lines indicate mean over 10 runs). Left: target $\mathcal{N}(4,2)$. Right: target mixture Model3 (mixture Pareto–Normal).
D.3 Moment-Agnostic Optimal Transport via Monotonicity-Penalized ISL
Our ISL framework applies to any probability law, including heavy-tailed distributions lacking finite higher-order moments. Indeed, the rank statistic in Eq. 1 is always well-defined, whereas the classical $p$–Wasserstein distance is only finite when both distributions possess finite $p$-th order moments.
Unique transport in one dimension.
In one dimension, any continuous map that pushes a simple base law $p_{z}$ onto a target $p$ must coincide with one of two inverses of the base cdf. If $F_{z}$ and $F$ denote the cdfs of $p_{z}$ and $p$, then
$f_{+}(z)=F^{-1}\bigl{(}F_{z}(z)\bigr{)},\quad f_{-}(z)=F^{-1}\bigl{(}1-F_{z}(z)\bigr{)}.$
The monotone map $f_{+}$ is in fact the unique optimal transport in $\mathbb{R}$. To recover this map, we augment our ISL loss with a monotonicity penalty.
Monotonicity-constrained training.
Given a batch of inputs $\{x_{i}\}_{i=1}^{N}$ sorted as
$x_{(1)}\leq x_{(2)}\leq\cdots\leq x_{(N)},\qquad f_{\theta}(x_{(i)})\;=\;\text{model output at }x_{(i)},$
we define
$\mathrm{Penalty}\;=\;\frac{1}{N}\sum_{i=1}^{N-1}\max\bigl{\{}0,\;f_{\theta}(x_{(i)})-f_{\theta}(x_{(i+1)})\bigr{\}},$
which is zero if and only if $f_{\theta}$ is non-decreasing. The overall training objective becomes
$\mathcal{L}(p,\tilde{p})\;=\;d_{K}(p,\tilde{p})\;+\;\lambda\,\frac{1}{N}\sum_{i=1}^{N-1}\max\bigl{\{}0,\;f_{\theta}(x_{(i)})-f_{\theta}(x_{(i+1)})\bigr{\}},$
where $d_{K}(p,\tilde{p})$ is our rank-based discrepancy and $\lambda>0$ weights the monotonicity constraint. As $\lambda\to\infty$, any violation of monotonicity incurs infinite cost, forcing $f_{\theta}$ to converge to the unique optimal transport map $f_{+}$.
This loss not only recovers the optimal transport in one dimension under minimal smoothness, but also extends beyond the Wasserstein framework to handle distributions with heavy tails.
We evaluate three training objectives—Dual-ISL with monotonicity penalty, 1-Wasserstein, and 2-Wasserstein—on a suite of heavy-tailed target distributions using a five-layer MLP with ELU activations and layer widths $[16,16,32,32,16,1]$. Each model is trained for $10^{4}$ epochs via vanilla gradient descent with a fixed learning rate of $10^{-2}$; we avoid adaptive optimizers to ensure that observed differences stem solely from the loss functions.
Performance is measured by two complementary metrics.
  1. 1.
    Kolmogorov–Smirnov distance (KSD): the maximum absolute deviation between the empirical cdfs of the real pdf $p$ and that of the generated distribution $\tilde{p}$.
  2. 2.
    Tail-fit error $A_{\mathrm{CCDF}}$: the area between the log–log complementary cdfs of real and generated samples, defined for $n$ data points by
    $A_{\mathrm{CCDF}}=\sum_{i=1}^{n}\Bigl{[}\log\bigl{(}F_{p}^{-1}(i/n)\bigr{)}-\log\bigl{(}\tilde{F}_{\tilde{p}}^{-1}(i/n)\bigr{)}\Bigr{]}\,\log\!\bigl{(}\tfrac{i+1}{i}\bigr{)},$
    where $F_{p}^{-1}$ and $\tilde{F}_{\tilde{p}}^{-1}$ are the inverse empirical ccdfs of $p$ and $\tilde{p}$, respectively.
Figure 8 compares the Dual-ISL loss (with monotonicity penalty) against the 1-Wasserstein loss on a Pareto–Normal mixture. Figure (a) shows that Dual-ISL accurately recovers the true transport map despite the heavy Pareto tails. Figure (b) demonstrates that the 1-Wasserstein loss fails to learn the correct mapping under heavy-tailed behavior. Figure (c) plots both losses over 10 000 training epochs: the Wasserstein loss oscillates and does not converge, whereas the Dual-ISL loss decreases smoothly and reliably, highlighting its stability and robustness.
(a) Dual-ISL
(b) Wasserstein
(c) Optimal-Transport
Figure 8: Comparison of transport objectives on a Pareto–Normal mixture: (a) Dual-ISL with monotonicity penalty, (b) 1-Wasserstein, and (c) Training dynamics over 10000 epochs, plotting Dual-ISL loss (solid) and 1-Wasserstein loss (dashed). The Dual-ISL curve decreases smoothly and converges reliably, whereas the Wasserstein loss oscillates and does not settle.
Table 3 shows that Dual-ISL (with monotonicity penalty) not only matches but frequently outperforms the classical OT baselines (1-Wasserstein and 2-Wasserstein) for $K=10$. For the moderately heavy-tailed $\mathrm{Cauchy}(5,10)$, 1-Wasserstein attains the lowest KSD, yet Dual-ISL achieves a smaller $A_{\mathrm{CCDF}}$, indicating superior tail alignment. As the tail heaviness grows in $\mathrm{Cauchy}(5,20)$ and $\mathrm{Pareto}(1,1)$, Dual-ISL outperforms both OT metrics on both measures—note that 2-Wasserstein is undefined for Pareto due to its infinite second moment—underscoring Dual-ISL’s robustness where classical OT either diverges or loses precision. Finally, on the multimodal $\mathrm{Model}_{3}$, Dual-ISL yields the lowest KSD and $A_{\mathrm{CCDF}}$, demonstrating that its rank-statistic formulation reliably recovers the unique monotone optimal-transport map in one dimension while avoiding the instability and non-differentiability of conventional OT losses.
Target Dual-ISL 1-Wasserstein 2-Wasserstein KSD ACCDF KSD ACCDF KSD ACCDF $\mathrm{Cauchy}(5,10)$ $0.069\pm 0.059$ $\mathbf{19.401\pm 0.539}$ $\mathbf{0.037\pm 0.024}$ $21.437\pm 0.429$ $0.504\pm 0.267$ $28.678\pm 7.018$ $\mathrm{Cauchy}(5,20)$ $\mathbf{0.045\pm 0.010}$ $\mathbf{17.201\pm 0.423}$ $0.046\pm 0.014$ $21.123\pm 1.290$ $0.668\pm 0.168$ $42.398\pm 11.578$ $\mathrm{Pareto}(1,1)$ $\mathbf{0.120\pm 0.053}$ $\mathbf{21.769\pm 0.072}$ $0.240\pm 0.064$ $23.676\pm 0.072$ $0.950\pm 0.158$ $\mathrm{Model}_{3}$ $\mathbf{0.131\pm 0.017}$ $\mathbf{19.406\pm 1.303}$ $0.148\pm 0.097$ $31.561\pm 2.769$ $0.219\pm 0.046$ $22.031\pm 1.387$
Table 3: KSD and ACCDF (mean $\pm$ std) for Dual-ISL (with monotonicity penalty) vs. Wasserstein baselines at $K=10$.
D.4 Empirical proof of the convergence rate
Our aim is to empirically validate Equation 3. To do this, we train a neural network using the same architecture as in our previous experiment. The network is fed an input pdf $\mathcal{N}(0,1)$ and is tasked with approximating a target pdf defined as a mixture of Cauchy distributions. We estimate $\tilde{p}$ via a kernel density estimator and compute the second derivative of $q$ using central finite differences with a sixth-order expansion. Each experiment is repeated $10$ times, and the mean results are shown in Figures 9.
(a) Fixed $K=10$. NN trained with $N=1000$ samples, lr $10^{-3}$.
(b) Varying $K$. Each run uses 1000 epochs, $N=1000$, lr $10^{-3}$.
Figure 9: Empirical convergence of dual-ISL’s Bernstein approximation (cf. Eq. 3). The solid blue curve shows the mean theoretical upper bound $\|q_{K}-1\|_{\infty}\leq(K+1)^{3}d_{K}$, and the dashed red curve shows the observed $\|q-1\|_{\infty}$.
D.5 Density estimation
D.5.1 1D density estimation
We employ the same fully-connected NN and training hyperparameters as in Appendix D.1. Once training converges, we approximate the implicit density with Equation 5. We restate the latter here for convenience
$p_{K}(x)$ $=\widehat{\tilde{p}}(x)\sum_{m=0}^{K}\mathbb{Q}_{K}(m)\,\tilde{b}_{m,K}\bigl{(}\widehat{\tilde{F}}(x)\bigr{)}\,,$
where
$\mathbb{Q}_{K}(m)\;=\;\Pr\bigl{(}A_{K}=m\bigr{)},\quad\tilde{b}_{m,K}(u)=\binom{K}{m}u^{m}(1-u)^{\,K-m}\,.$
The computation proceeds in three steps
  1. 1.
    Monte Carlo estimation of weights
    For each $m=0,\dots,K$, estimate
    $\mathbb{Q}_{K}(m)=\Pr\bigl{(}\#\{\,x_{i}\leq x\}=m\bigr{)}$
    by sampling $K$ independent latent vectors $z_{i}\sim\mathcal{N}(0,I)$, computing $x_{i}=f_{\theta}(z_{i})$, and counting how many satisfy $x_{i}\leq x$. Repeat $M$ times and take empirical frequencies:
    ^$\widehat{\mathbb{Q}}_{K}(m)=\frac{1}{M}\sum_{j=1}^{M}\mathbf{1}\Bigl{(}\#\{x_{i}^{(j)}\leq x\}=m\Bigr{)}.$
  2. 2.
    Empirical cdf estimation
    Draw $N$ samples $\{x_{i}\}_{i=1}^{N}$ from the trained generator, sort them in ascending order, and form the empirical cdf
    $\widehat{\tilde{F}}(x)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{x_{i}\leq x\}.$
  3. 3.
    Finite-difference density
    Approximate the density of the push-forward distribution by a first-order finite difference:
    $\widehat{\tilde{p}}(x)\approx\frac{\widehat{\tilde{F}}(x+\Delta)-\widehat{\tilde{F}}(x)}{\Delta},\qquad\Delta\ll 1.$
We evaluate Dual-ISL on six univariate target distributions. For each target, we estimate the mixture weights ^$\widehat{\mathbb{Q}}_{K}$ via $10^{5}$ Monte Carlo trials, compute the empirical cdf $\widehat{\tilde{F}}(x)$ from $10^{5}$ samples drawn from the trained generator, and form the density estimate using a first-order finite difference with $\Delta=0.1$. Table 4 reports the average Kolmogorov–Smirnov distance (over 10 independent runs) for the estimated density using Equation 5 with $K\in\{2,5,10\}$ versus a Gaussian kernel density estimator using the same number of samples and Silverman’s rule for bandwidth selection. Figure 10 illustrates density estimates for three representative targets—true density (solid red), ISL estimate (dashed blue), and KDE (dotted green)—all plotted on common axes to facilitate direct comparison of bias and tail behavior. As shown, Dual-ISL achieves the best results for every target except the Gaussian case, in which KDE with a Gaussian kernel performs marginally better.
Target Dual-ISL (K=2) Dual-ISL (K=5) Dual-ISL (K=10) KDE
$\mathcal{N}(4,2)$ $0.0202$ $0.0178$ $0.0167$ $\mathbf{0.0110}$
Cauchy$(1,2)$ $0.0237$ $0.0253$ $\mathbf{0.0184}$ $0.2013$
Pareto(1,1)11(1,1)( 1 , 1 ) $0.0302$ $\mathbf{0.0203}$ $0.0252$ $0.3872$
Mixture1 $0.0395$ $\mathbf{0.0095}$ $0.0120$ $0.0156$
Mixture2 $0.0171$ $0.0167$ $\mathbf{0.0070}$ $0.0145$
Mixture3 $0.1853$ $0.1786$ $\mathbf{0.0741}$ $0.1644$
Table 4: Mean Kolmogorov–Smirnov distance (over 10 runs) for Dual-ISL versus Gaussian KDE. We used ^$\widehat{\mathbb{Q}}_{K}$ estimated with $10{,}000$ trials, the empirical CDF $\widehat{\tilde{F}}$ from $10{,}000$ samples, and a finite-difference step $\Delta=0.1$.
(a) Cauchy
(b) Model2
(c) Model3
Figure 10: Density estimates for three univariate distributions using the ISL estimator (blue dashed line) and Gaussian kernel density estimator (KDE; green dotted line). Figure (a) shows the true Cauchy target density, while Figures (b) and (c) display the corresponding ISL and KDE estimates for Model2 and Model3, respectively. All panels share identical axes to facilitate direct comparison of estimator bias and tail behavior.
D.5.2 2D density estimation
To generalize our implicit estimator to data in $\mathbb{R}^{d}$, let $\{s_{\ell}\}_{\ell=1}^{m}\subset\mathbb{S}^{d}$ be random unit vectors. Denote by
$(s\#\hat{p})(y)\;=\;\text{ISL-estimate of the 1D pdf at }y=s^{\top}x$
the push-forward density along $s$. Then for any query point $x\in\mathbb{R}^{d}$,
$\hat{p}(x)$ $=\frac{1}{m}\sum_{\ell=1}^{m}\bigl{(}s_{\ell}\#\hat{p}\bigr{)}\bigl{(}s_{\ell}^{\top}x\bigr{)}.$
This Monte Carlo slicing—averaging one-dimensional ISL estimates—yields a consistent multivariate density approximation without ever constructing a full $d$-dimensional kernel.
In Figure 11, we compare the two-dimensional density estimates produced by our sliced Dual-ISL method against a Gaussian kernel density estimator with Silverman’s rule for bandwidth selection. Across all experiments, the generator is a four-layer MLP that maps 2D standard normal noise to the data space, with each hidden layer comprising 32 $\tanh$ units. We train for 100 epochs using 1000 samples per epoch, and approximate the sliced discrepancy by averaging over $m=10$ random projections with $K=10$. For the KDE baseline, we use 10000 samples to construct each density estimate.
ISL KDE
Dual Moon
Refer to caption Refer to caption
Figure 11: Two-dimensional density estimates on synthetic Dual Moon dataset. Left column: sliced Dual-ISL (averaged over $m=10$ random projections with $K=10$); right column: Gaussian KDE using Silverman’s bandwidth rule. The generator is a four-layer MLP with 32 units per hidden layer, trained for 100 epochs with 1000 samples per epoch; the KDE baseline uses 10000 samples.
D.6 Experiments on 2D distributions
We begin by evaluating three synthetic 2D benchmarks, each with a distinct topology:
  1. 1.
    Dual Moon: A bimodal Gaussian mixture (two disconnected modes).
  2. 2.
    Circle Gaussian: An eight-component Gaussian mixture arranged in a circle.
  3. 3.
    Two Ring: A “double-ring” distribution consisting of two concentric circular supports.
Our aim is to test whether the sliced dual-ISL method can fully recover these supports, including disconnected or non-convex regions. We compare against both normalizing flows and GANs, measuring performance by
  • KL divergence between the learned and true densities, and
  • Visual coverage of the true support via sample plots.
In all experiments—GAN, WGAN, and dual-ISL- the generator was a four-layer MLP that maps two-dimensional standard normal noise into data-space samples, with each hidden layer comprising 32 units and tanh activations. For the GAN and WGAN variants, the discriminator (or critic) adopted a similar four-layer MLP but with 128-unit hidden layers using ReLU activations and a final sigmoid output. Every model was trained for 1 000 epochs with a batch size of 1000 under the Adam optimizer. In the adversarial setups, we swept the critic-to-generator update ratio from 1:1 to 5:1 and chose the learning rate from $\{10^{-2},10^{-3},10^{-4},10^{-5}\}$. For ISL, we fix $K=10$, drew $N=1000$ samples per projection, averaged over $L=10$ random projections, and used a constant learning rate of $10^{-3}$.
For the normalizing flow baseline we adopted the RealNVP architecture of [dinh2016density], consisting of four affine-coupling layers whose scale and translation nets are two-hidden-layer MLPs with 32 ReLU units each. RealNVP was trained under the same 1 000-epoch, batch-size-1 000 protocol, but with a fixed learning rate of $5\times 10^{-5}$ as in [Stimper2023].
Figure 12 exposes the fundamental limitations of popular generative models on complex, multimodal data. GANs frequently collapse to a subset of modes, omitting entire regions of the true support. Normalizing flows, by enforcing invertibility, preserve the topology of the base distribution but struggle to represent disconnected clusters, instead “bridging” them with thin density filaments. In contrast, sliced Dual-ISL accurately recovers each connected component of the support—exemplified by the clean separation of the two moons—while still covering the full data manifold. On the Circle-of-Gaussians task, however, a low projection order $K$ can allow leakage between rings; raising $K$ sharpens the estimate and eliminates this spillover. A hybrid strategy (“Dual-ISL + GAN”), in which a model is first trained with Dual-ISL for 100 epochs and then fine-tuned adversarially, combines the stability and full-support coverage of Dual-ISL with the precision of a GAN, yielding the most faithful reconstructions. Table 5 reports KL divergences between each model and the true distribution, confirming that both Dual-ISL and especially the Dual-ISL + GAN variant outperform all baselines.
Target Real NVP WGAN ISL dual-ISL dual-ISL+GAN
Dual Moon
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Circle Gauss
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Two Rings
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Comparison of various generative methods on three 2D toy distributions. Each row (left) shows the target, and each column the outputs of Real NVP, WGAN, ISL-slicing, Dual-ISL, and dual-ISL(pretrained)+GAN, respectively.
Dataset Real NVP GAN WGAN ISL dual-ISL dual-ISL+GAN
Dual Moon 1.77 1.23 1.02 0.43 0.35 0.21
Circle of Gaussians 2.59 2.24 2.38 1.61 0.44 0.35
Two Rings 2.69 1.46 2.74 0.56 0.43 0.29
Table 5: KL-divergence (lower is better) of different generative models on 2D toy benchmarks.
Figure 13 overlays the true data samples (scatter points) with Gaussian-KDE contours computed from $10^{5}$ points generated by sliced dual-ISL. These contours closely match the underlying support of both the Dual Moon and Two Rings datasets, accurately delineating disconnected modes and concentric rings. This visualization confirms that sliced dual-ISL not only recovers the global topology of complex 2D manifolds but also captures fine geometric details in the learned density.
(a) Dual Moon
(b) Two Rings
Figure 13: Scatter plots of the true Dual Moon (left) and Two Rings (right) datasets, overlaid with density contours learned by Sliced dual-ISL and approximated via Gaussian KDE on $10^{5}$ generated samples. Models were trained for 1000 epochs with $L=20$ random projections and a batch size of $N=2000$.
D.7 High dimensional experiments
In this section, we evaluate the slicing dual-ISL method on high-dimensional image generation. In particular, we incorporate the dual-ISL objective into the training of a DCGAN generator [radford2015unsupervised] and benchmark the resulting architecture on the MNIST and Fashion-MNIST datasets. We report precision (a proxy for sample fidelity) and recall (a proxy for sample diversity) to assess the quality and diversity of the generated images [sajjadi2018assessing]. Our models were trained for 40 epochs with a batch size of 128, except for the pretrained models, which were first pretrained with sliced dual-ISL for 20 epochs (using 20 random projections) and then trained as a DCGAN for 40 epochs.
In Table 6, we report our results alongside those of other implicit generative models. On MNIST, our simple ISL-based model achieves recall rates comparable to much more complex, multi-discriminator GANs [durugkar2016generative, choi2022mcl] (5 discriminator each), despite using far fewer parameters. Furthermore, by pretraining the DCGAN generator with sliced dual-ISL and then fine-tuning under the standard adversarial loss, we attain state-of-the-art precision scores.
On Fashion MNIST, our model matches MCL-GAN in recall and—while we do not quite reach GMAN’s recall performance—our precision scores exceed theirs. This demonstrates that even with a simpler architecture, ISL can deliver competitive recall and precision results across diverse image-generation benchmarks.
Dataset Method F-score P&R
[…]
MNIST (28×28) dual-ISL (m=20) $85.00$ p m 0.32 $95.17$ p m 1.76 $84.85$ p m 1.20 $95.35$ p m 1.39
dual-ISL (m=50) $85.69$ p m 0.29 $95.81$ p m 1.24 $85.55$ p m 1.11 $96.23$ p m 1.98
DCGAN $93.58$ p m 0.64 $75.66$ p m 1.46 $93.85$ p m 1.45 $75.43$ p m 2.56
dual-ISL + DCGAN $93.58$ p m 0.84 $95.82$ p m 1.61 $94.03$ p m 1.82 $96.68$ p m 2.42
GMAN $97.60$ p m 0.70 $96.81$ p m 1.71 $97.60$ p m 1.82 $96.80$ p m 2.42
MCL-GAN $\mathbf{97.71\pm 0.19}$ $\mathbf{98.49\pm 1.57}$ $\mathbf{97.70\pm 1.33}$ $\mathbf{98.50\pm 2.15}$
FMNIST (28×28) dual-ISL (m=20) $81.84$ p m 0.11 $91.08$ p m 1.83 $81.48$ p m 1.43 $91.49$ p m 2.15
dual-ISL (m=50) $83.90$ p m 0.09 $91.18$ p m 1.57 $84.08$ p m 1.31 $92.92$ p m 1.23
DCGAN $86.14$ p m 0.11 $88.92$ p m 1.51 $86.60$ p m 1.58 $88.97$ p m 1.33
dual-ISL + DCGAN $91.43$ p m 0.19 $91.87$ p m 1.57 $91.88$ p m 1.35 $92.42$ p m 1.47
GMAN $90.97$ p m 0.09 $\mathbf{95.43\pm 1.12}$ $90.90$ p m 1.33 $\mathbf{95.50\pm 2.25}$
MCL-GAN $\mathbf{97.62\pm 0.09}$ $92.97$ p m 1.28 $\mathbf{97.70\pm 1.33}$ $92.90$ p m 2.31
Table 6: Quantitative comparison of generative models on MNIST and Fashion-MNIST ($28\times 28$) using F1/8, F8 (weighted harmonic mean of Precision and Recall, with $\beta=1/8$ and $\beta=8$ respectively), Precision, and Recall (mean ± std) in %. Results are shown for dual-ISL with $m=20$ and $m=50$ random projections, standard and Wasserstein DCGAN variants (with and without ISL pretraining), GMAN, and MCL-GAN. Boldface highlights the best score in each column per dataset. Higher is better.
Figure 14 shows the class-frequency distributions obtained by our sliced dual-ISL model (40 epochs, $m=50$ random projections) versus a standard DCGAN. Our model produces all ten digit classes in nearly uniform proportions—closely matching the true uniform distribution—while the DCGAN exhibits pronounced class imbalance. To compute these frequencies, we generated 10 000 samples from each model and classified them with a pretrained digit recognizer. In a Kolmogorov–Smirnov test for uniformity (on 10 000 samples), the sliced dual-ISL model achieved $p=0.070$, compared to $p=0.642$ for DCGAN, indicating a significantly better match to the ideal uniform distribution.
Finally, Figure 15 presents an ablation study of precision and recall in dual-ISL on MNIST, comparing the effects of using 20 versus 100 random projections. Even with only 20 projections, the model already achieves strong performance, and increasing the number of projections yields only marginal gains in these quality metrics.
00111122223333444455556666777788889999000.10.10.10.10.20.20.20.2Frecuency00111122223333444455556666777788889999000.10.10.10.10.20.20.20.2
Figure 14: Class-frequency distributions of MNIST samples generated by two different models. Left: dual-ISL; Right: standard DCGAN. Each bar represents the proportion of generated images assigned to each digit class (0–9), illustrating that dual-ISL produces a more uniform coverage across all classes, whereas the DCGAN exhibits notable biases toward certain digits. The dashed line indicates the ideal uniform distribution across all classes.
Figure 15: Precision (left) and recall (right) on the MNIST test set after 50 training epochs, using $m=20$ and $m=100$ random projections.
(a) DCGAN
(b) DCGAN pretrained with ISL
Figure 16: Figure 16(a) shows digits generated by a standard DCGAN, while Figure 16(b) shows samples from a DCGAN pretrained with ISL. The ISL-pretrained model exhibits significantly greater diversity across all classes, whereas the vanilla DCGAN produces an overabundance of ‘1’s.
(a) DCGAN
(b) DCGAN pretrained with ISL
Figure 17: Comparison of Fashion-MNIST samples: 17(a) generated by DCGAN trained 40 epochs, and 17(b) generated by a DCGAN pretrained with dual-ISL. The ISL-pretrained model demonstrates greater class diversity and improved precision.
Appendix E Pseudocodes
Below we summarize two variants of the dual-ISL training procedure. Algorithm 1 presents the basic likelihood-free, rank-based update in the one-dimensional case. Algorithm 2 extends this to multi-dimensional data by drawing random one-dimensional projections and averaging the resulting rank losses.
Algorithm 1 dual-ISL Training
0:  Generator network $f_{\theta}$, real samples $\{y_{i}\}_{i=1}^{N}$, batch size $M$, rank draws $K$, epochs $T$, learning rate $\eta$
0:  Trained parameters $\theta$
1:  for epoch $=1$ to $T$ do
2:     for each minibatch $B\subset\{y_{i}\}_{i=1}^{N}$ of size $M$ do
3:        Sample noise $z\sim p_{z}$
4:        Generate fictious sample $\tilde{y}=f_{\theta}(z)$
5:        Initialize histogram $\mathbf{q}\leftarrow\mathbf{0}$
6:        for $t=1,\dots,\lfloor M/K\rfloor$ do
7:           $\{y_{i}\}_{i=1}^{K}\leftarrow\text{RandomSubset}(B,\,K)$draw $K$ real samples from the minibatch}
8:           $a_{K}\leftarrow\sum_{i=1}^{K}\mathrm{SoftIndicator}[\tilde{y}\leq y_{i}]$differentiable count of the $A_{K}$ statistic}
9:           $\mathbf{q}\leftarrow\mathbf{q}+\mathrm{SoftHotEncoding}(a_{K},\ \text{length}=K+1)$accumulate a differentiable one-hot into the histogram}
10:        end for
11:        $\mathbf{q}\leftarrow\mathrm{normalize}(\mathbf{q})$
[…]
14:     end for
15:  end for
16:  return  $\theta$
Algorithm 2 Dual-ISL with Random Projections
0:  Generator $f_{\theta}$, real data $\{y_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$, batch size $M$, rank draws $K$, projection draws $L$, epochs $T$, learning rate $\eta$
0:  Learned parameters $\theta$
1:  for epoch = 1 to $T$ do
2:     for each minibatch $B\subset\{y_{i}\}$ of size $M$ do
3:        Sample noise $z\sim p_{z}$
4:        Generate fictious sample $\tilde{y}=f_{\theta}(z)$
5:        Initialize histogram $\mathbf{q}\leftarrow\mathbf{0}$
6:        for $\ell=1$ to $L$ do
7:           Sample random unit vector $v_{\ell}\sim\mathrm{Uniform}(S^{d-1})$
8:           for $t=1,\dots,\lfloor M/K\rfloor$ do
9:              $\{y_{i}\}_{i=1}^{K}\leftarrow\text{RandomSubset}(B,\,K)$draw $K$ real samples from the minibatch}
10:              Compute projections $\tilde{u}=v_{\ell}^{\top}\tilde{y}$ and $u_{i}=v_{\ell}^{\top}y_{i}$ for all $i$
11:              $a_{K}\leftarrow\sum_{i=1}^{K}\mathrm{SoftIndicator}[\tilde{u}\leq u_{i}]$differentiable count of the $A_{K}$ statistic}
12:              $\mathbf{q}\leftarrow\mathbf{q}+\mathrm{SoftHotEncoding}(a_{K},\ \text{length}=K+1)$accumulate a differentiable one-hot into the histogram}
13:           end for
14:           $\mathbf{q}\leftarrow\mathrm{normalize}(\mathbf{q})$
15:           $\mathrm{loss}\leftarrow\|\,\mathbf{q}-\tfrac{1}{K+1}\mathbf{1}_{K+1}\|_{1}$
16:        end for
17:        Compute loss: $\mathrm{projection\_loss=\mathrm{mean}(loss)}$
18:        Update: $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\,\mathrm{projection\_loss}$
19:     end for
20:  end for
21:  return  $\theta$
Appendix F Experimental Setup
All experiments were performed on a MacBook Pro running macOS 13.2.1, equipped with an Apple M1 Pro CPU and 16 GB of RAM. When GPU acceleration was required, we used a single NVIDIA TITAN Xp with 12 GB of VRAM. Detailed hyperparameter settings for each experiment are provided in the corresponding sections. An anonymous repository containing all code and data is available at https://anonymous.4open.science/r/dual-isl-6633. The code will also be included in the supplementary materials in a folder.
Appendix G Limitations
While our invariant statistical loss (ISL) framework eliminates the need for adversarial critics and guarantees strong convergence, it faces a critical trade-off when extended to high-dimensional data via ISL-slicing. Specifically, it requires $m$ random projections: a large $m$ enhances fidelity but incurs steep computational costs, whereas a small $m$ may overlook key anisotropic features. To address this, future research should design adaptive strategies for choosing or weighting projections—potentially drawing on recent advances in slicing-Wasserstein theory—to maximize information gain per projection. Moreover, exploring alternative projection methods (such as data-dependent or learned mappings) and establishing rigorous convergence bounds that link $m$ to both convergence rate and approximation error will be essential for fully automating and optimizing ISL’s performance.
By viewing ISL as the “cost” of rearranging generated samples to match real data points, we uncover its direct relationship with optimal transport. In essence, the permutation that sorts one sample set against another defines an explicit coupling—much like the Monge map—between the model and data distributions. Future work should formalize this correspondence, harnessing optimal transport tools to both analyze ISL’s theoretical properties and develop faster, more principled algorithms.
Appendix H Potential Societal Impact
Implicit generative models unlock powerful data-synthesis capabilities but also pose dual-use risks. Although our ISL framework could produce highly realistic outputs—such as photorealistic faces or authentic-sounding voices—it could likewise be misused to generate deep fakes. At the same time, we have shown that ISL excels at modeling heavy-tailed distributions, which helps ensure rare or minority subpopulations are neither over- nor under-represented. Furthermore, the closed-form density estimation inherent in ISL offers a transparent window into what is otherwise a black-box process, improving explainability and enabling practitioners to audit and adjust model behavior for fairness. By combining high-fidelity synthesis with built-in safeguards and interpretability, ISL paves the way for more responsible, equitable deployment of generative technologies.